### v1

# MLVU: A Comprehensive Benchmark for Multi-Task Long Video Understanding

Abstract
The evaluation of Long Video Understanding (LVU) performance poses an important but challenging research problem. Despite previous efforts, the existing video understanding benchmarks are severely constrained by several issues, especially the insufficient lengths of videos, a lack of diversity in video types and evaluation tasks, and the inappropriateness for evaluating LVU performances. To address the above problems, we propose a new benchmark, called MLVU [^1] MLVU Repository: https://github.com/FlagOpen/FlagEmbedding. (Multi-task Long Video Understanding Benchmark), for the comprehensive and in-depth evaluation of LVU. MLVU presents the following critical values: 1) The substantial and flexible extension of video lengths, which enables the benchmark to evaluate LVU performance across a wide range of durations. 2) The inclusion of various video genres, e.g., movies, surveillance footage, egocentric videos, cartoons, game videos, etc., which reflects the models’ LVU performances in different scenarios. 3) The development of diversified evaluation tasks, which enables a comprehensive examination of MLLMs’ key abilities in long-video understanding. The empirical study with 20 latest MLLMs reveals significant room for improvement in today’s technique, as all existing methods struggle with most of the evaluation tasks and exhibit severe performance degradation when handling longer videos. Additionally, it suggests that factors such as context length, image-understanding quality, and the choice of LLM backbone can play critical roles in future advancements. We anticipate that MLVU will advance the research of long video understanding by providing a comprehensive and in-depth analysis of MLLMs.
1 Introduction
Large language models (LLMs) are growing into a general solution for numerous AI tasks [6; 39]. In recent years, it becomes increasingly emphasized to extend LLMs with multi-modal capabilities and thus bring the Multi-modal LLM, namely, MLLM. Remarkably, it has been made possible for today’s MLLMs to perceive information in texts, images, videos, etc., and solve complicated problems in physical environments [1; 38]. Along with the development of MLLMs, new benchmarks are continuously created to facilitate comprehensive and in-depth analysis of MLLMs [50; 27; 9; 21].
However, it remains a great challenge to evaluate the MLLMs’ long-video understanding (LVU) performances given the following limitations. Firstly, the majority of existing video understanding benchmarks are made up of short videos [47; 21; 19; 31; 15], whose lengths can be merely a few seconds. As a result, they are insufficient to reflect the MLLMs’ long-video understanding capabilities. Secondly, there is a severe lack of diversity for the existing LVU benchmarks in terms of video genres and evaluation tasks. On one hand, the existing benchmarks are usually created based on one type of videos, e.g., movies [35; 22]. On the other hand, it may only contain one specific type of task in each individual benchmark, for example, the captioning task in [47], the temporal perception task in [49], and the action understanding task in [41; 11]. Due to the above limitations, existing benchmarks have difficulty in performing a comprehensive evaluation of LVU performances. Last but not least, many previous evaluation tasks are not properly designed for LVU, as they can be solved without using the complex information from long videos. For example, many questions are simply about one single frame in the long videos [35]. Besides, numerous others are about popular movies and celebrities [22], which can be answered directly by MLLMs based on the textual prompts.
Conceptually, the MLLMs are expected to handle any type of long video and accomplish any type of related tasks. Therefore, the evaluation of LVU should emphasize two important properties: length and diversity. Based on such a principle, we propose a novel benchmark called MLVU (Mult-task Long Video Understanding Benchmark), which presents the following critical advantages.
  • It makes a substantial extension for the video length. MLVU is created based on long videos of diversified lengths, ranging from 3 minutes to 2 hours. The average video length is about 12 minutes, which makes it much longer than most of the existing benchmarks. Additionally, each video is further segmented so that evaluation tasks can be created w.r.t. different video clips (e.g., summarization for the first 3 minutes, the first 6 minutes, and the entire duration of the video). Therefore, it is able to flexibly evaluate the MLLMs’ performance across different video lengths.
  • It covers a wide variety of video genres. On one hand, it contains different types of real-world videos, such as movies, documentaries, surveillance videos, and ego-centric videos. On the other hand, it also includes typical simulated videos, like games and cartoons. As a result, it is able to comprehensively reflect the MLLMs’ performance across different application scenarios.
  • It introduces diversified evaluation tasks tailored for LVU. There are 9 different tasks in MLVU, which jointly examine a wide range of MLLMs’ key abilities, such as reasoning, captioning, recognition, summarization, etc. Both multi-choice and free-form generation tasks are included in MLVU, which reflect the MLLMs’ performances in handling different forms of tasks. Finally, some of the tasks are designed to leverage global information from entire videos, while others need to utilize proper local information from specific clips. Therefore, it can provide a unified perspective for the degree of completeness and nuance in understanding long videos.
We extensively investigate 20 popular MLLMs with MLVU, which brings in several critical insights. Firstly, long-video understanding remains a technically challenging problem for the existing MLLMs. While GPT-4o [^1] https://openai.com/index/hello-gpt-4o/ achieves the leading performance in the experiment, it only attains an average score of 64.6% in multi-choice tasks. All methods struggle with tasks requiring fine-grained information from entire videos, such as action counting, ordering, and summarization. Additionally, all methods experience significant performance drops as video lengths increase. Secondly, a significant gap persists between open-source and proprietary models. Although some of the latest open-source MLLMs can handle very long videos and excel at single-image tasks, they notably lag behind GPT-4o in understanding long videos. Finally, the empirical results underscore the influential factors in LVU, such as the extension of context length, the improvement of image understanding ability, and the utilization of strong LLM-backbones. In addition to the benchmark’s overall conclusion, the individual tasks enable the fine-grained analysis of MLLMs’ performances in each of the specialized aspects. Therefore, we anticipate the benchmark to assist in improving MLLMs’ long-video understanding capabilities by providing insights into their current strengths and weaknesses.
2 Related Work
Multimodal Large Langugae Models.
Multimodal large language models (MLLMs) have gained wide interests from both academic and industrial sectors. Recent studies have made significant progresses in this field by integrating LLM backbones with visual encoders and adapters, and fine-tuning the entire architecture through visual instruction [25; 55; 7]. Based on the same philosophy, MLLMs have been further developed for video processing using video instruction datasets and specialized video adapters [52; 28; 20; 48; 23; 21]. However, most existing models are optimized for short videos, typically under one minute, due to the difficulty in establishing sufficient context for longer videos. To address this challenge, researchers have explored compact representations of videos. For example, LLaMa-Vid [22] compresses each video frame into two tokens, allowing the model to handle videos several hours long. Methods like MovieChat [35] and MA-LMM [12] introduce specialized memory components for recursive video processing. Additionally, it is also explored to make selective usage of frames or clips from long videos based on retrievers or agents [46; 33; 40]. Despite these progresses, it remains an open problem for MLLMs to effectively handle long videos.
Video Understanding Benchmarks.
With the unprecedented interest in MLLMs, the creation of benchmarks for these models has become increasingly emphasized (as advanced by MMMU [50], MME [9], and many other pioneering works). In video understanding, the research community has made significant efforts as well, particularly for short videos. There are specialized benchmarks for temporal perception [49; 42], action understanding [42; 41], video classification [14], video reasoning [45; 44], and video captioning [47; 30]. Recently, MVBench [21] provides a comprehensive short-video benchmark to evaluate general capabilities via question-answering. For long video understanding, people seek to leverage long-form videos, like movies, to create benchmarks. For example, the LVU dataset [43] presents movie understanding tasks, such as predicting release years and identifying character relationships. Similarly, LLaMA-Vid [22] created a movie question-answering dataset based on MovieNet [13]. Despite using long videos, many questions are about common-sense knowledge of the movies. As a result, they could be directly answered without using the video’s information. In contrast, MovieChat [35] makes use of diversified videos and avoids specific character names or plot details in its questions. Considering that many of the questions target on exact time segments, it potentially degrades the tasks to short-video or image understanding problems. Beyond movies, there are other task-specific benchmarks like EgoSchema [29], which presents video reasoning tasks using first-person footage from Ego4D [11]. However, these specialized benchmarks only focus on one single aspect of MLLMs, rather than offering a comprehensive analysis of long video understanding. Therefore, it remains essential to develop a comprehensive benchmark with carefully designed tasks to effectively evaluate MLLMs’ capabilities in understanding long videos.
3 MLVU: Multi-task Long Video Understanding Benchmark
In this section, we start with an overview of MLVU, which highlights its constitution and explains its values over the previous works. Then, we discuss how each evaluation task is constructed in MLVU.
3.1 Overview
MLVU is a comprehensive benchmark made up of 2593 evaluation tasks [^2] Only consider the tasks for the entire videos. Others for fine-grained video segments are not accounted. (shown as the right side of Figure 1), which belong to 9 task categories tailored for long-video understanding. These categories include 1) Topic Reasoning, 2) Anomaly Recognition, 3) Video Summarization, 4) Needle Question Answering, 5) Ego Reasoning, 6) Plot Question Answering, 7) Sub-Scene Captioning, 8) Action Count, and 9) Action Order. The benchmark is distinguished by the following features.
Figure 1: Statistical overview of our LVBench dataset. Left: Distribution of Video Duration; Middle: Distribution of Source Types for Long Videos; Right: Quantification of Each Task Type.
Substantial extension of video length.
MLVU is made up of videos of diversified lengths, spanning from 3 min to more than 2 hours (Figure 1 Left). The average length is about 12 min, which is much longer than the existing LVU benchmarks. Besides, each video is further partitioned as incremental segments, e.g., the first 3 min, the first 6 min, and the entire video, where tasks are created for each individual segment. Thus, the MLLMs can be flexibly evaluated across different video lengths.
Diversified video categories.
MLVU offers a comprehensive collection of videos of different categories (Figure 1 Middle). There are typical real-world videos, which includes movies [51; 35], documentaries [35], TV series, egocentric videos [11], and surveillance footage [37]. Meanwhile, there are also important simulated videos from animated series and game videos [8].
Diversified evaluation tasks.
MLVU also provides a diversified collection of evaluation tasks, which are closely related to the common visual capabilities of MLLMs, such as reasoning, captioning, recognition, perception, and summarization (Figure 1 Right). All the tasks are tailored for LVU. That is to say, the tasks need to be solved based on the in-depth understanding of video. Some of tasks are to examine whether the global information from the entire video can be effectively utilized (holistic LVU); while others focus on whether the MLLMs can make precise usage of proper local information within the long video (detail LVU). Additionally, both multi-choice and free-form generation tasks are included in MLVU, which help to examine MLLMs’ capabilities in handling different task formats.
Figure 2: Examples of MLVU. There are nine tasks designed to evaluate the holistic, single-detail, and multi-detail LVU capabilities of MLLMs. The MLLMs are asked to solve the problem (with the ground-truth answers marked in blue) based on the long video input and textual prompt.
3.2 Construction of MLVU
The evaluation tasks of MLVU which can be categorized into three types: 1) holistic LVU, which needs to be solved by making use of the global information from the entire video; 2) single-detail LVU, which needs to leverage one critical plot within the long video; 3) multi-detail LVU, which calls for the joint utilization of multiple plots within the long video. The construction process of MLVU is discussed w.r.t. the above three categories. To facilitate the discussion, we define ULVC (Universal Long Video Collection) as the universal collection of miscellaneous long videos, which includes movies, documentaries, TV-series, ego-centric videos, surveillance footages, animated series, game videos, et al. (more details about ULVC is presented in the Appendix B).
3.2.1 Holistic LVU
Topic Reasoning (TR). The topic reasoning task requires MLLMs to respond to questions about the principal subject of a long video, as shown with Figure 2 (a). This includes elements such as the video’s genre, pivotal events, or primary settings. Videos for TR tasks are sourced from movies, documentaries, animated series, egocentric videos, and game videos from ULVC. All questions and answers undergo manual annotation [^3] Detailed information and annotation guidelines for annotators are presented in Appendix E, resulting in a total of 264 questions. TR tasks are formatted as multiple-choice questions, with the model’s performance assessed based on accuracy.
Anomaly Recognition (AR). The anomaly recognition task involves identifying the anomalous behavior within a surveillance footage (Figure 2 b). We leverage the surveillance video clips from UCF Crime dataset [37] for this task. The selected video clips are longer than three minutes. We create 200 questions based on the original annotations provided by the dataset. The AR task is also conducted in the multiple-choice format, whose performance is measured by accuracy.
Video Summarization (VS). This task requires MLLMs to summarize the key events in a long video (Figure 2 c). We select the narrative-rich videos from ULVC for this task, including movies, TV series, documentaries, and animated series. There are 217 selected videos in total, whose summaries are manually annotated. During evaluation, the MLLMs are prompted with "Please summarize the main content of this video". We employ GPT-4 to assess the generated summaries by comparing with the annotation results. Details about annotation and evaluation are presented in Appendix E.3 and F.3.
3.2.2 Single-Detail LVU
Needle Question-Answering (NQA). Needle-In-the-Haystack-Search (NIHS) is a popular evaluation task for long-context LLM [26]. Taking the inspiration from NIHS, we create Needle Question-Answering (NQA), shown as Figure 2 (d). In this task, the MLLM is required to answer a question related to a specific segment (referred as needle) within a long video (referred as background video). The needles are short video clips sampled from WebVid [5], while the background videos are sampled from our ULVC. The needle is randomly inserted into the background video, where a question-answer pair is annotated. By incorporating necessary details, the question can always correspond to the needle without ambiguity. During evaluation, the MLLM needs to infer the location of the needle based on the details provided in the question, and solve the problem on top of the needle’s information. The NQA task is structured as multiple-choice, whose performance is measured by accuracy.
Ego Reasoning (ER). Ego-centric videos capture a series of consecutive actions from a first-person perspective. The MLLM needs to reason for a question about a specific behavior in the video, e.g., predicting for the event which is correlated or satisfies a certain causal relationship with the behavior (Figure 2 e). Both videos and QA annotations are collected from the NLQ task of Ego4D [11]. The ER task is structured as multiple-choice, with a total of 352 questions created for this task.
Plot Question-Answering (PQA). In this task, the MLLM needs to reason for questions about a plot in a narrative video, shown as Figure 2 (f). The video is sampled from the movies, TV series, and animated series in our ULVC. There are 539 question-answer pairs created by manual annotation. During annotation, the human annotators are asked to only provide necessary details about the plot but not to suggest any objective hints, e.g., the two characters in the example video are referred as cat and mouse, rather than Tom and Jerry. Therefore, it can prevent the question from being short-cut by the MLLM’s common-sense knowledge (more details about PQA can be found in the Appendix E.6).
Sub-Scene Captioning (SSC). In this task, the MLLM needs to generate the caption for a sub-scene in a long video. The long videos in SSC are sampled from the Movie101 dataset [51], while the questions and answers are manually annotated. During annotation, the human annotator is asked to provide a detailed description for the sub-scene as the ground-truth answer. Besides, they need to offer necessary clues in their questions such that the referred sub-scenes can be identified without ambiguity. During evaluation, we employ GPT-4 [1] to measure the quality of caption in comparison with the ground-truth. Details about annotation and evaluation are presented in Appendix E.7 and F.3.
3.2.3 Multi-Detail LVU
Action Order (AO). In this task, the MLLM needs to predict the right order for a sequence of actions (Figure 2 h). The actions are presented by short video clips, called probes. The probes are formulated in two different ways. One is made up of clips from the Kinetics dataset [14], where each clip represents a distinct action. The other one is from the consecutive clips of an action in the ActivityNet-Caption dataset [16]. The probes are inserted into a long background video, which is sampled from ULVC. There are 259 AO questions in total. The task is structured as a multiple-choice prblem, where the right order is selected from the misleading options provided by the annotator.
Action Count (AC). This task requires the MLLM to count the occurrences of an action within a long video (Figure 2 i). Each action corresponds to multiple short probe clips sampled from the Kinetics dataset [14]. The probes of an action are inserted into a long background video sampled from ULVC. We also perform manual examination to ensure that the inserted action does not exist in the original background video. A total of 206 evaluation instances have been created. The AC task is structured as a multiple-choice problem, with performance measured by accuracy.
4 Experiments and Analysis
4.1 Settings
We perform a comprehensive investigation of 20 MLLMs based on MLVU, both open-source and proprietary. The experimental MLLMs can be partitioned into three categories. 1. Image MLLMs, which are primarily fine-tuned by image-related instructions. We consider the following models for this category: Otter-I [18], LLaVA-1.6 [24], InternVL-1.5 [7], Claude3-Opus [2], Qwen-VL-Max [4], and GPT-4 Turbo [1]). 2. Short Video MLLMs, which are fine-tuned by short-video related instructions. This category includes: Otter-V [18], mPlug-Owl-V [48], Video LLaMA-2 [52], Video ChatGPT [28], VideoChat [20], VideoChat2 [21], and Video-LLaVA [23]. 3. Long Video MLLMs, which are optimized for their long-video understanding capability. This category includes: MiniGPT4-Video [3], LLaMA-VID [22], Movie-LLM [36], MA-LMM [12], MovieChat [35], TimeChat [34], and GPT-4o [32]. For Image MLLMs, we leverage their multi-image inference capabilities to process segmented frames from original videos. In the case of Video MLLMs, we employ either a uniform sampling strategy or a frame rate sampling strategy for video processing. All models are evaluated based on their official implementations or available APIs, where the evaluation is conducted in a zero-shot manner. More details about the evaluation are provided in the Appendix F.
4.2 Overall Performance
The overall evaluation results for all investigated MLLMs are shown in Table 1. The individual performance is reported for each individual task; meanwhile, the average performances are reported for the multiple-choice (M-Avg) and generation tasks (G-Avg), respectively. We can obtain two primary observations from the demonstrated results.
On one hand, the recently released GPT-4o [32] achieves an overwhelming advantage in our benchmark. It outperforms the rest of methods by a big margin in terms of the average performances, with an M-Avg of 64.6% (within 0-100%) and a G-Avg of 5.80 (within 0.0-10.0). Besides, it also maintains the leading position in every individual task. As for the open-source models, InternVL-1.5 [7] achieves the highest average performance in multiple-choice tasks, whose M-Avg scores reaches 50.4%. Such a performance is impressive knowing that it even slightly goes beyond GPT-4 Turbo [1]. Meanwhile, LLaMA-VID [22] presents the top performance for the generation tasks, whose G-Avg score reaches 4.22. However, such a score is significantly lagging behind GPT-4 Turbo and GPT-4o.
On the other hand, although GPT-4o achieves a huge advantage in comparison with other methods, it actually struggles to handle most of the tasks in the benchmark. For example, it only ends up with 64.8% in dealing with the needle question-answering (NQA) task. Whereas its analogue tasks in the text domain, e.g., NIHS (Needle-In-the-HayStack-Search) and Passkey Retrieval, can be effectively conquered by many of the existing long LLMs [10; 53]. At the same time, it exhibits even less reliability given tasks like ego-reasoning (ER), action ordering (AO), and action count (AC), while other baseline methods produce even worse performances in these scenarios. The above observations indicate that long-video understanding remains a tough challenge for today’s MLLMs.
In addition to the primary conclusions from the overall performances, we can also make the following interesting observations about the individual tasks. First of all, the multiple-choice holistic tasks, i.e., topic retrieval (TR) and anomaly recognition (AR), present much higher differentiation than other tasks. Proprietary MLLMs, like GPT-4o, GPT-4-turbo, and superior open-source models, like InternVL-1.5, can accurately solve such problems; meanwhile, many other popular MLLMs still fail to generate meaningful performances. Knowing that the two tasks only require an overall understanding of the long videos, they can serve as a preliminary indicator of MLLMs’ LVU ability.
Besides, it’s hard to deal with tasks which need nuanced understanding of multiple details. Although several MLLMs can handle single-detail LVU tasks to some extent, their performances suffer from catastrophic degradation when addressing multi-detail LVU tasks. Most methods, except for GPT-4o, fail entirely in action order (AO) and action count (AC) tasks. Additionally, most approaches struggle with the summarization task, which require recalling multiple nuanced details from long videos.
Finally, the existing long-videos MLLMs (except for GPT-4o) are no better than those primarily trained from short-video or image related instructions. Although such models can intake much longer videos (i.e. much more frames than other baselines), it turns out that the extended input information is not effectively utilized. On the contrary, Video-LLaVA can achieve a relatively competitive performance, especially in multiple-choice, despite that it can only intake 8 frames.
As a brief conclusion, although today’s MLLMs can deal with some preliminary LVU tasks, it remains a tough challenge to achieve an in-depth understanding of nuanced information within long videos.
Methods Input Holistic Single Detail Multi Detail M-Avg G-Avg
TR AR VS NQA ER PQA SSC AO AC
Full mark - 100 100 10 100 100 100 100 10 100 100 10
Image MLLMs
[Uncaptioned image] Otter-I [18] 16 frm 25.0 25.0 2.18 25.1 25.0 24.9 4.12 13.1 25.2 23.3 3.15
[Uncaptioned image]LLaVA-1.6 [25] 16 frm 60.6 41.0 2.11 43.1 38.4 41.0 4.35 25.5 25.7 39.3 3.23
[Uncaptioned image]InternVL-1.5 [7] 16 frm 78.8 67.0 3.16 52.7 43.5 54.4 4.88 32.8 23.8 50.4 4.02
[Uncaptioned image]Claude-3-Opus [2] 16 frm 67.2 43.5 3.11 21.6 40.2 47.8 3.66 18.2 16.7 36.5 3.39
[Uncaptioned image]Qwen-VL-Max [4] 16 frm 67.4 63.5 2.71 40.3 40.9 43.3 5.21 25.0 14.8 42.2 3.96
[Uncaptioned image] GPT-4 Turbo [1] 16 frm 79.5 68.0 4.49 45.9 47.4 60.6 6.21 26.5 16.1 49.2 5.35
Short Video MLLMs
[Uncaptioned image]Otter-V [18] 16 frm 24.6 26.0 2.38 28.2 27.6 22.3 4.23 15.1 26.7 24.4 3.31
[Uncaptioned image]mPLUG-Owl-V [48] 16 frm 28.0 25.0 2.36 24.5 31.8 27.3 5.31 21.2 23.3 25.9 3.84
[Uncaptioned image]VideoChat [20] 16 frm 33.0 32.0 2.31 27.0 32.1 27.6 5.01 24.3 28.6 29.2 3.66
[Uncaptioned image]Video-ChatGPT [28] 100 frm 26.9 24.0 2.31 40.3 42.0 29.9 5.48 25.1 31.1 31.3 3.90
[Uncaptioned image]Video-LLaMA-2 [52] 16 frm 54.5 41.5 2.34 39.4 33.5 35.4 5.22 18.5 25.7 35.5 3.78
[Uncaptioned image]VideoChat2 [21] 16 frm 74.6 51.5 2.57 42.0 47.4 43.8 5.04 22.8 29.6 44.5 3.81
[Uncaptioned image]Video-LLaVA [23] 8 frm 71.6 57.0 2.43 53.2 45.2 48.4 5.25 20.1 35.9 47.3 3.84
Long Video MLLMs
[Uncaptioned image]MovieChat [35] 2048 frm 29.5 25.0 2.33 24.2 24.7 25.8 3.23 28.6 22.8 25.8 2.78
[Uncaptioned image]Movie-LLM [36] 1 fps 30.0 29.0 2.88 29.6 24.7 24.1 5.00 20.5 24.8 26.1 3.94
[Uncaptioned image]TimeChat [34] 96 frm 23.1 27.0 2.54 24.5 28.4 25.8 4.29 24.7 32.0 30.9 3.42
[Uncaptioned image]LLaMA-VID [22] 1 fps 50.8 34.5 3.22 30.1 32.7 32.5 5.22 23.9 27.8 33.2 4.22
[Uncaptioned image]MA-LMM [12] 1000 frm 51.9 35.5 2.12 43.1 38.9 35.8 4.80 25.1 24.3 36.4 3.46
[Uncaptioned image] MiniGPT4-Video [3] 90 frm 70.9 52.5 2.64 49.0 48.6 44.5 4.07 23.2 23.0 44.5 3.36
[Uncaptioned image]GPT-4o [32] 0.5 fps 87.4 74.5 4.90 64.8 57.1 65.1 6.69 56.7 46.3 64.6 5.80
[…]
Figure 3: Experimental performances on MLVU under various video lengths: GPT4-turbo (orange), MiniGPT4-Video (red), VideoChat2 (green), Video-LLaMA-2 (blue), and MA-LMM (cyan).
4.3 Detailed Analysis
We analyze the impact from video length and three factors: context length, image understanding (IU) ability, LLM-backbone. All the factors are empirically critical to MLLMs’ LVU performances.
In the first place, we evaluate MLLMs’ performances across various video lengths. For this purpose, we introduce a derivative dataset alongside MLVU, called MLVU Time-ladder. In this dataset, the same kinds of evaluation tasks are created for videos of variant lengths, including 180s, 360s, and 600s (more details presented in Appendix D). As shown in Figure 3, the performances of all models tend to decline as the video length grows, which indicates that the existing MLLMs’ LVU abilities are severely constrained by the video length. By comparison, image models, like GPT-4-turbo [1], and short-video models, like VideoChat2 [21], VideoLLaMA [52], are more vulnerable to the growth of video length, while long video models, like MiniGPT4-Video, can be relatively more resilient.
     Impact of Context Length Impact of IU Impact of LLM
  Model Context Len. M-Avg Model MMMU (Val) M-Avg Model LLM M-Avg
  MGV 16 41.1 Otter 32.2 24.4 VLM2 Vicuna-7B 24.3
90 44.5$\uparrow$3.4 LLaVA 35.8 39.3$\uparrow$14.9 Vicuna-13B 35.5$\uparrow$11.2
GPT-4o 16 54.2 GPT4V 58.1 49.2 MGV LLAMA-7B 33.7
256 64.6$\uparrow$10.4 GPT-4o 63.8 54.2$\uparrow$5.0 Mistral-7B 44.5 $\uparrow$10.8
Table 2: Detailed analysis about the impact from context length, image understanding (IU) ability, and LLM Backbone. MGV: MiniGPT4-Video, GPT4V: GPT-4-turbo, VLM2: Video-LLaMA-2.
We further examine MLLMs’ performances (measured by M-Avg) under varying context lengths. Specifically, we increase MiniGPT4-Video’s input from 16 to 90 frames and GPT-4o’s input from 16 to 256 frames (the left side of Table 2). As the input length extends, both models consistently show improved performance. To investigate the impact of MLLMs’ image understanding (IU) ability, we reference the experiment results from MMMU [50] (the middle of Table 2). It’s evident that the MLLMs’ LVU performances basically align with their image understanding performances in MMMU. Finally, we compare MLLMs with different backbones (the right side of Table 2). The results show that LVU performances improve with larger (Vicuna-13B vs. Vicuna-7B) and more capable (Mistral-7B vs. Llama-2-7B) backbone encoders. These observations indicate that LVU is the result of multiple complex factors, with the ability to perceive longer videos and effectively utilize the perceived information being crucial for the improvement of LVU.
5 Conclusion & Discussion
This paper presents MLVU, a novel benchmark for the assessment of long video understanding. With several critical innovations: the substantial extension of video lengths, the inclusion of various video genres, and the development of diversified LVU-oriented evaluation tasks, the new benchmark is able provide a comprehensive and in-depth analysis for MLLMs’ long-video understanding performance. The empirical study on MLVU reveals LVU remains a technically challenging problem for today’s state-of-the-art MLLMs. Future advancements may call for the joint optimization of complex factors, such as context length, image understanding ability, and even LLM backbones. We anticipate this benchmark will facilitate future research in long-video understanding of MLLMs.
Limitations. While MLVU basically covers the major dimensions of long-video understanding, it can still be improved for even better comprehensiveness. For instance, there can be tasks related to high-resolution videos, or more specific tasks such as tracking and low-level processing. Our work will be a persistent effort, where new tasks and video types will be continually introduced.
Appendix A Overview of Appendix
  • B: Collecting Details of our Universal Long Video Collection (ULVC).
  • C: Distribution of Video Durations for Each Task.
  • D: Details of the MLVU Time-Ladder
  • E: Annotation Details of MLVU.
  • F: Details of Baselines and the Evaluation Process.
  • G: Explorations of Video Retrieval Augmented Generation.
  • H: More Experimental Results.
  • I: More Visualized Examples of MLVU.
  • J: Licensing, Hosting and Maintenance Plan.
  • K: Datasheet.
  • Checklist
Appendix B Collecting Details of our Universal Long Video Collection (ULVC)
In the initial stage of our Multi-task Long Video Understanding (MLVU) benchmark creation, we first collected long-form videos from a variety of sources to form our Universal Long Video Collection (ULVC). The entirety of the long videos incorporated into our MLVU benchmark were selected, edited, or synthesized from ULVC.
Specifically, our ULVC consists of a diverse set of 757 long videos. This collection includes 168 movies sourced from the Movie101 dataset [51] and the MovieChat dataset [35], as well as 60 documentaries from the MovieChat dataset [35]. Additionally, it comprises 65 game videos from the MineDojo dataset [8], 200 surveillance videos from the UCF-Crime dataset [37], and 100 ego-centric videos from the Ego4D dataset [11]. Furthermore, our team independently collected 72 cartoons and 92 TV series to enrich the dataset.
It’s important to clarify that the quantity of videos in the ULVC does not directly correspond to the number of videos and questions in our MLVU benchmark, which are 1334 and 2593 respectively. For example, a two-hour movie from the ULVC might be utilized in its entirety for the Sub-Scene Captioning task, or it could be segmented into several approximately 10-minute clips for the Video Summary task, or even used as a background video for synthetic video generation. Moreover, a single video could be annotated with multiple questions simultaneously.
Appendix C Distribution of Video Durations for Each Task
Figure 4 illustrates the detailed video duration distribution for each task in our MLVU, featuring videos of varying lengths from three minutes to over 120 minutes.
Figure 4: Video duration distributions of different tasks in our MLVU.
Appendix D Details of the MLVU Time-Ladder
As discussed in Section 3.1, most tasks in our MLVU are subject to segment-level annotation. This approach provides us with the flexibility to adjust the length of the video without requiring additional human annotators. Building on this strategy, as mentioned in Section 4.3, we have generated a derivative dataset, MLVU Time-Ladder, which includes videos of varying durations - specifically 3, 6, and 10 minutes. This dataset allows us to investigate how video duration impacts LVU task difficulty.
Specifically, during the annotation process of the VS task, we guided annotators to delineate the summarization in accordance with the initial 3 and 6-minute segments. For the PQA and SSC tasks, we requested annotators to identify the segments within the extended video where the pertinent answers are located. In the case of the ego reasoning task, the Ego4D dataset [11] already comprises the intervals where the answers reside. Lastly, for the synthetic tasks of NQA, AO, and AC, we possess the capability to directly generate the necessary video lengths.
Appendix E Annotation Details of MLVU
E.1 Topic Reasoning (TR).
The questions and corresponding answers for the TR task were meticulously annotated by human annotators, following the specific guidelines illustrated in Figure 5. We required the annotators to design questions related to the reasoning of the video topic, rather than focusing on the creation of questions about minor details. More visualized examples of TR task can be found in Figure 14.
Figure 5: Annotation Guidelines for the Topic Reasoning Task.
E.2 Anomaly Recognition (AR).
The anomaly recognition task did not involve manual annotation. We utilized videos exceeding three minutes in duration, extracted from the UCF-Crime dataset [37]. We also modified the original labels to fit a multiple-choice format.
E.3 Video Summarization (VS).
The ground truth data for the VS task were derived from manual annotations. We instructed the annotators to use pronouns instead of specific character names in all annotations. This guideline stemmed from the inherent constraints of most existing MLLMs, which generally lacked the capacity to process audio or subtitles. This made it difficult for these models to identify specific characters. The annotation instructions and examples provided to the annotators are elaborated in Figure 6. More visualized examples of VS task can be found in Figure 14.
Figure 6: Annotation Guidelines for the Video Summarization Task.
E.4 Needle Question-Answering (NQA).
We leveraged the GPT-4 [1] and the detailed video caption data from the WebVid dataset [5] to facilitate a semi-automated generation of annotated questions and answers for the NQA task. Initially, we selected video clips from WebVid, which we refered to as needle clips. The corresponding captions of these needle clips were then fed into GPT-4, which generated question-answer pairs based on the information encapsulated in the captions. The specific prompt provided to GPT-4 is depicted in Figure 7. The generated questions were carefully crafted to focus on a particular detail within the needle clip. These questions were structured to incorporate the maximum number of hints to effectively guide MLLMs in grounding the content of the needle within the context of the longer video. Following this, we randomly selected longer background videos from our ULVC and manually ensured that the scene indicated by the needle’s question did not feature in these background videos. The final step involves integrating the needle into the longer video, thereby producing the final needle question video. More visualized examples of NQA task can be found in Figure 15.
Figure 7: The prompt provided to GPT-4 in the process of creating the question-answer pair for the Needle Question-Answering task.
E.5 Ego Reasoning (ER).
The video resources, questions, and correct responses used in the ER task were derived from the Natural Language Queries (NLQ) task within the Ego4D dataset [11]. This data was restructured to fit a multiple-choice question format.
E.6 Plot Question-Answering (PQA).
The PQA task’s questions and answers were annotated by human annotators, following specific guidelines illustrated in Figure 8. We instructed the annotators to craft questions that probe into the intricate plot details encapsulated within the videos. These questions were designed to encompass both perception and reasoning aspects. We stipulated that both questions and their corresponding answers should avoid the use of specific character names or any objective hints, and should instead utilize pronouns. This approach was strategized to prevent potential information leakage, given that MLLMs often demonstrate a familiarity with the storylines of well-known movies and TV series. Such common-sense knowledge could potentially allow the MLLMs to answer questions correctly without the essential requirement of analyzing the input video.
Nonetheless, the complexity of character interactions and actions in longer videos poses a challenge to conveying plot details using only pronouns and feature descriptions. Previous datasets for plot question answering that avoided the use of character names often resulted in compromised question diversity and tended towards generalized queries. We illustrate this through a comparative analysis of TVQA [17], Moviechat [35], and our PQA dataset’s question word clouds in Figure 9. While TVQA provides a diverse range of questions, it does so by employing specific character names. In contrast, Moviechat avoids character names, but its questions are frequently overly broad, lack specific plot details, and exhibit diminished diversity. Our PQA dataset successfully navigates these challenges, offering a diverse range of questions without resorting to the use of character names. More visualized examples of PQA task can be found in Figure 15.
Figure 8: Annotation Guidelines for the Plot Question-Answering Task.
Figure 9: Word Cloud Comparison of questions in TVQA test set (left), MovieChat test set (middle), and our PQA (right). Notably, TVQA’s character-specific names require LLMs to recognize characters, risking reliance on pre-existing knowledge. In contrast, MovieChat questions are less diverse. Our PQA addresses these issues, providing enhanced usability and reliability.
E.7 Sub-Scene Captioning (SSC).
In the development process of the SSC task, we employed human annotators to generate both prompts and standard caption data. The specific guidelines provided to annotators are illustrated in Figure 10. Initially, the annotators identified a specific, easily referable sub-scene within a lengthy movie. Subsequently, they crafted a prompt replete with adequate clues to reference this scene, ensuring the uniqueness of these clues throughout the entire film. To prevent any leakage of information, the prompt was designed to exclude any character-specific names or objective hints, instead incorporating rich descriptive details to allude to the plot. Following this, the annotators produced a detailed caption for this sub-scene, and deconstructed the caption into multiple, non-redundant "scoring points" to facilitate quantitative assessment (the details of the evaluation metric can be found in Section F.3). More visualized examples of PQA task can be found in Figure 16.
Figure 10: Annotation Guidelines for the Sub-Scene Captioning Task.
E.8 Action Order (AO).
The videos, questions, and answers for the action order task were all synthetically generated. In order to maintain the high quality of our evaluation data, we adopted a dual-strategy approach. Firstly, we selected actions for the probe videos that were not commonly seen in most films, such as making jewelry and water skiing. Secondly, in the selection of background videos, we conducted a cursory review of the video content to further ensure that the actions referenced in the questions were not present in the video. This rigorous methodology ensured the reliability of our data.
E.9 Action Count (AC).
The process of data acquisition and annotation for the action count task closely mirrored that of the action order task. All videos, questions, and answers were synthetically generated. We employed a strategy consistent with the action order task to ensure the validity and reliability of our evaluation data.
Appendix F Details of Baselines and the Evaluation Process
F.1 Baselines
In this section, we detail the primary baselines evaluated on our MLVU. For image-based MLLMs, most available models lack multi-image inference capabilities. Consequently, we select Otter-I, LLaVA-1.6, and InternVL, which offer official multi-image implementations. Additionally, we include three proprietary models—Claude-3-Opus, Qwen-VL-Max, and GPT-4 Turbo—that provide APIs for multi-image inference. For available models, we estimate the maximum input frames based on their maximum LLM context length. For Claude and Qwen, the supported maximum image numbers are approximately 20; therefore, we select 16 frames to ensure fair comparisons. Similarly, we choose 16 frames for GPT-4 Turbo to maintain consistency. In terms of video MLLMs, we adhere to the default settings for frame sampling strategies. For instance, VideoChat2 uniformly samples 16 frames, whereas LLaMA-Vid samples 1 frame per second. Specifically, the GPT-4o can support a maximum of approximately 500 images when each image’s resolution is set to 512$\times$512 pixels. Consequently, we opt for a sampling rate of 0.5 fps to accommodate the majority of our videos. Table 3 and Table 4 show the LLM-backbone as well as evaluated weight links used in each MLLM.
     Image MLLMs Short Video MLLMs Long Video MLLMs
  Model LLM Model LLM Model LLM
  Otter-I MPT-7B Otter-V LLaMA-7B Movie-LLM Vicuna-7B
LLaVA-1.6 Vicuna-7B mPLUG-Owl-V LLaMA-7B MovieChat Vicuna-7B
InternVL-1.5 Vicuna-13B VideoChat Vicuna-7B TimeChat LLaMA-2 7B
Claude-3-Opus Claude3 Video-ChatGPT Vicuna-7B LLaMA-VID Vicuna-7B
Qwen-VL-Max Qwen Video-LLaMA-2 Vicuna-13B MA-LMM Vicuna-7B
GPT-4 Turbo GPT-4 VideoChat2 Vicuna-7B MiniGPT4-Video Mistral-7B
Video-LLaVA Vicuna-7B GPT-4o GPT4
Table 3: LLM-Backbone in different MLLMs
     Image MLLMs Short Video MLLMs Long Video MLLMs
  Model Weight Model Weight Model Weight
  Otter-I Link Otter-V Link Movie-LLM Link
LLaVA-1.6 Link mPLUG-Owl-V Link MovieChat Link
InternVL-1.5 Link VideoChat Link TimeChat Link
Claude-3-Opus - Video-ChatGPT Link LLaMA-VID Link
Qwen-VL-Max - Video-LLaMA-2 Link MA-LMM Link
GPT-4 Turbo - VideoChat2 Link MiniGPT4-Video Link
Video-LLaVA Link GPT-4o -
Table 4: Link of evaluation weights. The specific version for “GPT-4 Turbo” and “GPT-4o” is “gpt-4-turbo-2024-04-09” and “gpt-4o-2024-05-13”.
F.2 Inference Detatils
We have developed two templates specifically for Multiple-Choice and Generation tasks, as illustrated in Figure 11. Distinct system prompts were designed to accommodate the differences between video-based and image-based MLLMs. Considering the variances in task requirements, we incorporated “option prediction guidance” into the Multiple-Choice template to aid in option extraction. Conversely, in Generation tasks, we do not implement any additional interventions but employ fixed-question guidance to enable models to respond to diverse task questions. In our evaluation, the templates are seamlessly integrated into the evaluation code of open-release models or available API of proprietary models.
F.3 Evaluation Metrics
For the evaluation of Multiple Choice tasks, we directly compute absolute accuracy by matching the predicted option with the ground truth. In Generation tasks, we develop multiple criteria for assessment and employ GPT-4 to rank the alignment between generated texts and the provided answers. As illustrated in Figure 12, we use “Accuracy” and “Relevance” to benchmark Sub-scene Captioning, and “Completeness” and “Reliability” to evaluate the capabilities of Video Summary.
Figure 11: Inference template for our MLVU.
Figure 12: Detailed prompt for evaluation of generation tasks in MLVU.
Appendix G Explorations of Video Retrieval Augmented Generation
As discussed in Section 4.3, most MLLMs are adversely affected by video length. Drawing inspiration from the use of Retrieval Augmented Generation (RAG) in video understanding, we have developed a zero-shot RAG strategy and seamlessly integrated it into existing MLLMs. Table 5 displays the performance comparison between the baseline models and the models employing our RAG strategy. It is noteworthy that all methods benefit from the RAG strategy in Needle QA, Ego Reasoning, and Plot QA. Conversely, minimal improvement is observed in Action Count, and a decrease is noted in Action Order and Overall Reasoning. This is primarily because RAG facilitates the retrieval of detail-oriented video clips, which makes models more likely to focus on answer-related cues in specific single-detail reasoning tasks. However, RAG exhibits limited capabilities in multi-detail reasoning and holistic understanding tasks, which require global perception and knowledge aggregation.
The pipeline of our video retrieval augmented generation is illustrated in Figure 13. Initially, a long video is uniformly divided into $N$ video clips, each containing $C$ frames. Subsequently, we employ a robust video feature extraction tool, LanguageBind [54] to extract clip embeddings $F_{I}\in\mathbb{R}^{N\times d}$, where $d$ represents the dimension of each clip embedding. We then compute the similarities between $F_{I}$ and the text embedding $F_{T}$, concatenating the top $K$ clips to enhance the model’s capability for question-answering. Given that many Video MLLMs are limited to processing only 16 frames, we have adjusted the settings for $C$ and $K$ to accommodate video retrieval in 16-second intervals. As discussed below, the RAG strategy excels in detail-oriented tasks but shows limitations in global understanding tasks. Moreover, it is relatively inefficient, requiring more than one minute to complete the process. Consequently, more effective approaches need to be developed for long video understanding tasks, and we aim to address this in future work.
Figure 13: Pipeline of our video retrieval augmented generation strategy.
  Model Settings Needle QA Ego Rea. Plot QA Action Or. Action Co. Anomaly Rec. Topic Rea.
  LLaVA-B - 43.1 38.4 41.0 25.5 25.7 41.0 60.6
  LLaVA-R C=2,K=8 50.7 45.7 49.7 26.3 26.7 40.8 59.8
C=4,K=4 53.5 43.5 50.6 25.9 29.6 39.9 58.5
C=8,K=2 55.2 42.6 50.3 25.1 30.1 40.6 59.5
  InternVL-B - 52.7 43.5 54.4 32.8 23.8 67.0 78.8
  InternVL-R C=2,K=8 77.2 52.6 61.4 30.1 36.4 57.9 69.2
C=4,K=4 76.3 51.4 59.9 29.3 36.9 58.3 69.4
C=8,K=2 77.8 48.9 61.6 31.7 33.0 60.2 62.3
  Video-LLaMA-B - 39.4 33.5 35.4 18.5 25.7 41.5 54.5
  Video-LLaMA-R C=2,K=8 61.4 42.6 38.8 17.4 17.5 35.7 48.5
C=4,K=4 58.9 42.6 39.1 17.8 23.8 36.0 49.3
C=8,K=2 62.0 38.4 36.2 25.5 18.0 38.5 51.0
  VideoChat2-B - 42.0 47.4 43.8 22.8 29.6 51.5 74.6
  VideoChat2-R C=2,K=8 72.1 53.7 55.5 21.6 30.1 45.8 68.2
C=4,K=4 72.4 55.4 53.4 22.4 31.1 45.3 68.9
C=8,K=2 73.8 53.1 55.3 22.0 31.6 46.6 69.7
  MiniGPT4-Video-B - 49.0 48.6 44.5 23.2 23.0 52.5 70.9
  MiniGPT4-Video-R C=2,K=8 60.6 44.3 47.4 23.2 23.7 42.8 60.9
C=4,K=4 60.3 44.6 46.9 26.3 23.8 42.6 60.7
C=8,K=2 56.3 44.6 46.6 27.4 24.8 45.0 47.5
Table 5: Quantitative results on video Retrieval Augmented Generation. “model-B” and “model-R” denote Baseline and RAG models respectively. We evaluate two image MLLMs and three video MLLMs in different settings.
Rank Model Score
  Refer to caption GPT-4o 87.4
Refer to caption GPT-4 Turbo 79.5
Refer to caption InternVL-1.5 78.8
4 VideoChat2 74.6
5 Video-LLaVA 71.6
6 MiniGPT4-Video 70.9
7 Qwen-VL-Max 67.4
8 Claude-3-Opus 67.2
9 LLaVA 60.6
10 Video-LLaMA-2 54.5
(a) Anomaly Recognition
Rank Model Score
  Refer to caption GPT-4o 74.5
Refer to caption GPT-4 Turbo 68.0
Refer to caption InternVL-1.5 67.0
4 Qwen-VL-Max 63.5
5 Video-LLaVA 57.0
6 MiniGPT4-Video 52.5
7 VideoChat2 51.5
8 Claude-3-Opus 43.5
9 Video-LLaMA-2 41.5
10 LLaVA 41.0
(b) Topic Reasoning
Rank Model Score
  Refer to caption GPT-4o 4.90
Refer to caption GPT-4 Turbo 4.49
Refer to caption LLaMA-VID 3.22
4 InternVL-1.5 3.16
5 Claude-3-Opus 3.11
6 Movie-LLM 2.88
7 Qwen-VL-Max 2.71
8 MiniGPT4-Video 2.64
9 VideoChat2 2.57
10 TimeChat 2.54
(c) Video Summarization
Rank Model Score
  Refer to caption GPT-4o 64.8
Refer to caption Video-LLaVA 53.2
Refer to caption InternVL-1.5 52.7
4 MiniGPT4-Video 49.0
5 GPT-4 Turbo 45.9
6 LLaVA 43.1
7 MA-LMM 43.1
8 VideoChat2 42.0
9 Qwen-VL-Max 40.3
10 Video-ChatGPT 40.3
(d) Needle QA
Rank Model Score
  Refer to caption GPT-4o 57.1
Refer to caption MiniGPT4-Video 48.6
Refer to caption GPT-4 Turbo 47.4
4 VideoChat2 47.4
5 Video-LLaVA 45.2
6 InternVL-1.5 43.5
7 Video-ChatGPT 42.0
8 Qwen-VL-Max 40.9
9 Claude-3-Opus 40.2
10 MA-LMM 38.9
(e) Ego Reasoning
Rank Model Score
  Refer to caption GPT-4o 65.1
Refer to caption GPT-4 Turbo 60.6
Refer to caption InternVL-1.5 54.4
4 Video-LLaVA 48.4
5 Claude-3-Opus 47.8
6 MiniGPT4-Video 44.5
7 VideoChat2 43.8
8 Qwen-VL-Max 43.3
9 LLaVA 41.0
10 MA-LMM 35.8
(f) Plot QA
Rank Model Score
  Refer to caption GPT-4o 6.69
Refer to caption GPT-4 Turbo 6.21
Refer to caption Video-ChatGPT 5.48
4 mPLUG-Owl-V 5.31
5 Video-LLaVA 5.25
6 Video-LLaMA-2 5.22
7 LLaMA-VID 5.22
8 Qwen-VL-Max 5.21
9 VideoChat2 5.04
10 VideoChat 5.01
(g) Sub-Scene Caption.
Rank Model Score
  Refer to caption GPT-4o 56.7
Refer to caption InternVL-1.5 32.8
Refer to caption MovieChat 28.6
4 GPT-4V 26.5
5 LLaVA 25.5
6 Video-ChatGPT 25.1
7 MA-LMM 25.1
8 Qwen-VL-Max 25.0
9 TimeChat 24.7
10 VideoChat 24.3
(h) Action Order
Rank Model Score
  Refer to caption GPT-4o 46.3
Refer to caption Video-LLaVA 35.9
Refer to caption TimeChat 32.0
4 Video-ChatGPT 31.1
5 VideoChat2 29.6
6 VideoChat 28.6
7 LLaMA-VID 27.8
8 Otter-V 26.7
9 LLaVA 25.7
10 Video-LLaMA-2 25.7
(i) Action Count
Rank Model Score
  Refer to caption GPT-4o 64.6
Refer to caption InternVL-1.5 50.4
Refer to caption GPT-4 Turbo 49.2
4 Video-LLaVA 47.3
5 MiniGPT4-Video 44.5
6 VideoChat2 44.5
7 Qwen-VL-Max 42.2
8 LLaVA-1.6 39.3
9 Claude-3-Opus 36.5
10 MA-LMM 36.4
(j) M-Avg
Rank Model Score
  Refer to caption GPT-4o 5.80
Refer to caption GPT-4 Turbo 5.35
Refer to caption LLaMA-VID 4.22
4 InternVL-1.5 4.02
5 Qwen-VL-Max 3.96
6 Video-ChatGPT 3.90
7 mPLUG-Owl-V 3.84
8 Video-LLaVA 3.84
9 VideoChat2 3.81
10 Video-LLaMA-2 3.78
(k) G-Avg
Table 6: Leaderboard of different tasks in MLVU.
(until 2024/06/01)
Appendix H More Experimental Results
H.1 Leaderboard of each task in MLVU
Table 6 shows the leaderboard of each task in our MLVU.
H.2 Comprehensive evaluation results on MLVU Time-ladder
Table 7, Table 8 and Table 9 show the experimental results on MLVU Time-ladder within 3 minutes, 6 minutes and 10 minutes. As anomaly recognition and topic reasoning are global tasks, thus, we can not annotate them in MLVU Time-ladder as the way provided in Appendix D.
  Model Needle QA Ego Rea. Plot QA Action Or. Action Co. Sub-Scene Cap. Video Sum. M-Avg
  Otter-I 25.0 25.0 25.4 25.0 25.0 4.04 2.59 25.1
LLaVA-1.6 48.0 43.0 45.2 29.5 28.5 4.48 2.40 38.8
InternVL-1.5 70.5 47.5 60.9 19.5 28.5 5.42 3.30 45.3
Claude-3-Opus 34.5 42.5 38.1 15.5 22.0 4.12 3.35 30.5
Qwen-VL-Max 49.5 42.2 50.3 27.7 16.3 5.58 2.95 37.2
Otter-V 30.5 32.0 18.3 21.0 29.5 4.32 2.63 26.3
mPLUG-Owl-V 31.0 29.5 22.3 29.0 24.0 5.35 2.68 27.2
Video-LLaMA-2 55.0 35.0 36.5 22.5 22.0 5.76 2.47 34.3
Video-ChatGPT 43.5 41.0 40.0 29.0 28.0 5.51 2.65 34.5
VideoChat 32.5 31.5 26.4 24.5 26.0 5.39 2.61 28.2
VideoChat2 54.0 53.5 50.3 30.0 31.0 5.11 3.25 43.8
Video-LLaVA 58.0 47.5 55.8 29.5 34.5 5.38 2.89 45.0
MiniGPT4-Video 57.0 44.5 47.2 27.5 26.5 4.38 2.98 40.5
LLaMA-VID 38.5 36.5 36.0 22.5 25.5 5.31 3.55 31.8
Movie-LLM 23.5 22.5 25.9 25.0 25.0 5.13 3.23 34.4
MA-LMM 48.0 43.0 41.6 27.5 23.5 4.55 2.25 36.7
MovieChat 32.5 29.5 27.4 21.0 21.0 3.60 2.60 26.3
TimeChat 30.5 30.5 31.0 27.5 24.5 4.29 2.92 28.8
Table 7: Quantitative results on MLVU Time-ladder (180s).“M-Avg” is calculated by the average of five multiple-choice tasks. Bold fonts mean the best results.
  Model Needle QA Ego Rea. Plot QA Action Or. Action Co. Sub-Scene Cap. Video Sum. M-Avg
  Otter-I 25.0 25.0 25.0 11.5 25.0 4.17 2.33 22.4
LLaVA-1.6 41.0 38.5 39.6 22.0 26.0 4.14 2.18 33.4
InternVL-1.5 66.5 48.5 58.4 34.0 24.0 5.23 3.12 46.2
Claude-3-Opus 32.0 44.7 42.6 30.2 15.5 3.75 3.13 33.0
Qwen-VL-Max 50.0 43.4 41.1 24.4 14.7 5.49 3.21 34.7
Otter-V 27.0 30.5 23.9 13.0 26.5 4.34 2.51 24.2
mPLUG-Owl-V 29.5 34.5 22.3 23.5 23.0 5.58 2.42 26.6
Video-LLaMA-2 43.0 36.0 40.1 23.0 28.0 5.62 2.45 34.0
Video-ChatGPT 42.5 35.0 27.4 23.5 32.0 5.22 2.56 32.1
VideoChat 28.5 27.5 19.8 25.5 30.0 5.32 2.38 26.3
VideoChat2 47.5 44.5 39.6 27.0 31.0 4.97 2.83 37.9
Video-LLaVA 55.5 47.5 47.7 22.0 35.0 5.25 2.62 41.5
MiniGPT4-Video 53.0 43.5 43.1 31.0 24.5 4.11 2.82 39.0
LLaMA-VID 30.5 39.0 28.4 19.5 30.5 5.39 3.29 29.6
Movie-LLM 26.0 22.5 32.0 22.5 26.0 5.16 3.14 25.8
MA-LMM 41.5 37.0 41.6 27.0 27.5 4.52 2.19 34.9
MovieChat 28.5 25.5 29.4 25.0 23.5 3.25 2.41 26.4
TimeChat 36.5 30.0 31.0 26.5 26.0 4.37 2.66 30.0
Table 8: Quantitative results on MLVU Time-ladder (360s).
  Model Needle QA Ego Rea. Plot QA Action Or. Action Co. Sub-Scene Cap. M-Avg.
  Otter-I 25.0 25.0 25.0 12.5 25.0 4.02 22.5
LLaVA-1.6 37.5 38.5 43.9 25.0 28.5 4.45 34.6
InternVL-1.5 56.0 46.0 55.6 33.0 25.0 5.17 43.1
Claude-3-Opus 27.0 38.5 39.8 35.7 15.0 3.88 31.2
Qwen-VL-Max 43.0 44.8 50.0 17.3 17.9 5.45 34.6
Otter-V 27.5 28.5 22.4 14.0 26.0 4.09 23.7
mPLUG-Owl-V 30.5 36.0 21.4 20.0 28.0 5.29 27.2
Video-LLaMA-2 39.0 39.0 29.6 21.5 25.5 5.62 30.9
Video-ChatGPT 41.0 40.5 28.6 24.5 32.5 5.37 33.4
VideoChat 25.5 34.5 26.0 28.0 30.0 4.91 28.8
VideoChat2 43.0 49.0 41.8 24.5 29.5 5.05 37.6
Video-LLaVA 54.0 49.0 52.0 21.0 29.5 5.21 41.1
MiniGPT4-Video 54.0 39.0 41.8 25.0 28.0 4.13 37.6
LLaMA-VID 32.0 36.5 34.2 27.0 27.0 5.39 31.3
Movie-LLM 26.5 32.5 27.0 21.5 24.5 5.25 26.4
MA-LMM 44.5 41.5 35.7 24.0 28.0 4.72 34.7
MovieChat 27.5 32.5 22.4 24.0 27.0 3.48 26.7
TimeChat 28.0 32.0 23.5 26.0 24.5 4.19 26.8
Table 9: Quantitative results on MLVU Time-ladder (600s).
Appendix I More Visualized Examples of MLVU.
Figure 14: More Examples of Topic Reasoning and Video Summarization Tasks.
Figure 15: More Examples of Needle Question Answering and Plot Question Answering Tasks.
Figure 16: More Examples of Sub-Scene Captioning.
Appendix J Licensing, Hosting and Maintenance Plan
Author Statement.
We bear all responsibilities for the licensing, distribution, and maintenance of our dataset.
License.
MLVU is under CC-BY-NC-SA 4.0 license.
Hosting.
MLVU can be viewed and downloaded on GitHub at https://github.com/FlagOpen/FlagEmbedding/tree/master/MLVU or on Huggingface at https://huggingface.co/datasets/MLVU/MVLU. We assure its long-term preservation for future reference and use. The annotations for questions and answers are provided in the JSON file format, while the raw videos are available in the MP4 format.
For the video files, we do not hold any copyright, and all video rights belong to the video authors. To facilitate user usage, under the premise of a user agreement that the data will be used solely for research purposes and not for commercial use, we provide the download methods for these videos. We have taken measures such as reducing resolution, modifying aspect ratio, and editing to minimize the impact on the original work rights. However, there is still a risk that copyright holders may request the removal of some data. If this happens, we will follow the practice of Movienet [13] to sparsely collect these video frames. This will not significantly impact our data usage, as all our annotations are focused on the visual information in the videos, unrelated to the audio. Moreover, most of the models we evaluate are essentially still frame-based video processing and do not involve audio processing. If retaining video frames is also not allowed, we will still preserve the annotation data and provide metadata for the corresponding videos.
Metadata.
Appendix K Datasheet
K.1 Motivation
For what purpose was the dataset created?
Answer: The creation of MLVU is to facilitate the evaluation and development of MLLM’s long video understanding capabilities. Compared to previous video understanding benchmarks, MLVU possesses a substantial extension of video length, diversified video categories, and diversified evaluation tasks. MLVU is the first comprehensive benchmark for long video understanding.
Who created the dataset (e.g., which team, research group) and on behalf of which entity (e.g., company, institution, organization)?
Answer: The MLVU is created by Junjie Zhou (BUPT, BAAI), Yan Shu (BAAI), Bo Zhao (BAAI), Boya Wu (BAAI), Shitao Xiao (BAAI), Xi Yang (BAAI), Yongping Xiong (BUPT), Bo Zhang (ZJU), Tiejun Huang (PKU, BAAI), and Zheng Liu (BAAI)
Who funded the creation of the dataset?
Answer: Beijing Academy of Artificial Intelligence, Beijing University of Posts and Telecommunications, Peking University, and Zhejiang University.
K.2 Composition
What do the instances that comprise the dataset represent? (e.g., documents, photos, people, countries)
Answer: Each instance in our dataset represents a long video ranging from 3 minutes to 2 hours in duration, a question, a standard answer, and for the multiple-choice task, there are also 4 options. Videos are stored in MP4 file format, while the questions, standard answers, and options are all stored in JSON format files.
How many instances are there in total (of each type, if appropriate)?
Answer: In total, we collect 2593 instances, including 2175 multiple-choice questions and 418 free-form generation questions. The number of distinct videos is 1334, as one video may correspond to multiple questions. The specific distribution of question types and video duration can be found in Figure 1 in the main paper.
Does the dataset contain all possible instances or is it a sample (not necessarily random) of instances from a larger set?
Answer: Only a part of the questions was sampled from existing datasets (the ego reasoning task and the anomaly recognition task). The rest of the questions were newly annotated by us (for example, plot question answering, sub-scene captioning, video summarization, topic reasoning), or they were substantially modified based on existing annotations (needle question answering, action count, action order). During the data collection and annotation process, we ensured that the instances within each task were diverse in terms of videos and questions.
Is there a label or target associated with each instance?
Answer: Yes, for the multiple-choice question type, each instance provides the correct option. For the free-form generation task, each instance comes with a standard caption or summary answer.
Is any information missing from individual instances?
Answer: All instances are complete.
Are relationships between individual instances made explicit (e.g., users’ movie ratings, social network links)?
Answer: Some instances may have the same video but different questions and answers. Each instance clearly indicates the corresponding video, and each video has a unique identifier.
Are there recommended data splits (e.g., training, development/validation, testing)?
Answer: No. Our MLVU is specifically designed for evaluation, with its core objective being to assess the capability of MLLMs to understand long-term video.
Are there any errors, sources of noise, or redundancies in the dataset?
Answer: No.
Is the dataset self-contained, or does it link to or otherwise rely on external resources (e.g., websites, tweets, other datasets)?
Answer: All data will be publicly accessible in the dataset repository. Our annotations will be stored in JSON format. As our collected videos include parts of movies, TV series, documentaries, and cartoons, we have taken measures such as reducing resolution, modifying aspect ratio, and editing to minimize the impact on the original work rights. However, there is still a risk that copyright holders may request removal of some data. If this happens, we will follow the practice of Movienet [13] to sparsely collect these video frames. This will not have much impact on our data usage, as all our annotations are focused on the visual information in the videos, unrelated to the audio. If retaining video frames is also not allowed, we will still keep the annotation data and provide metadata for the corresponding videos.
Does the dataset contain data that might be considered confidential?
Answer: No.
Does the dataset contain data that, if viewed directly, might be offensive, insulting, threatening, or might otherwise cause anxiety?
Answer: No.
K.3 Collection Process
The data collection process is described in Section 3.2 of the main paper and Appendix B, D, and E.
K.4 Uses
Has the dataset been used for any tasks already?
Answer: Yes, MLVU has been used to evaluate the long-video understanding capabilities of as many as 20 different MLLMs. For specific details, please refer to Section 4.
What (other) tasks could the dataset be used for?
Answer: MLVU is primarily used to comprehensively evaluate the long-video understanding capabilities of MLLMs. Since MLVU includes multiple different tasks, it can also be used to individually assess the ability of MLLMs or other video-specific models on a particular task, such as video summarization.
Is there a repository that links to any or all papers or systems that use the dataset?
Answer: No.
Is there anything about the composition of the dataset or the way it was collected and preprocessed/cleaned/labeled that might impact future uses?
Answer: As our collected videos include parts of movies, TV series, documentaries, and cartoons, we have taken measures such as reducing resolution, modifying aspect ratio, and editing to minimize the impact on the original work rights. However, there is still a risk that copyright holders may request removal of some data. If this happens, we will follow the practice of Movienet [13] to sparsely collect these video frames. This will not have much impact on our data usage, as all our annotations are focused on the visual information in the videos, unrelated to the audio. If retaining video frames is also not allowed, we will still keep the annotation data and provide metadata for the corresponding videos.
Are there tasks for which the dataset should not be used?
Answer: The MLVU cannot be used to evaluate the image or video generation capabilities of the MLLM.
K.5 Distribution
Will the dataset be distributed to third parties outside of the entity (e.g., company, institution, organization) on behalf of which the dataset was created?
Answer: Yes. The benchmark is publicly available on the Internet.
How will the dataset will be distributed (e.g., tarball on website, API, GitHub)?
Answer: The benchmark is available on GitHub at https://github.com/FlagOpen/FlagEmbedding/tree/master/MLVU or on Huggingface at https://huggingface.co/datasets/MLVU/MVLU.
Will the dataset be distributed under a copyright or other intellectual property (IP) license, and/or under applicable terms of use (ToU)?
Answer: CC-BY-NC-SA 4.0.
Have any third parties imposed IP-based or other restrictions on the data associated with the instances?
Answer: No.
Do any export controls or other regulatory restrictions apply to the dataset or to individual instances?
Answer: No.
K.6 Maintenance
Who will be supporting/hosting/maintaining the dataset?
Answer: The authors will be supporting, hosting, and maintaining the dataset.
How can the owner/curator/manager of the dataset be contacted (e.g., email address)?
Answer: Please contact the official email of our project (mlvubenchmark@gmail.com), or contact the one of the authors (zhoujunjie@bupt.edu.cn; shuyan9812@gmail.com)
Is there an erratum?
Answer: No. We will make announcements if there are any.
Will the dataset be updated (e.g., to correct labeling errors, add new instances, delete instances)?
If the dataset relates to people, are there applicable limits on the retention of the data associated with the instances (e.g., were individuals in question told that their data would be retained for a fixed period of time and then deleted)?
Answer: People may appear in the reference videos. People may contact us to exclude specific data instances if they appear in the reference videos.
Will older versions of the dataset continue to be supported/hosted/maintained?
If others want to extend/augment/build on/contribute to the dataset, is there a mechanism for them to do so?
Answer: If others wish to add data, they can apply to do so provided the data is compliant and reasonable. However, making other modifications based on our dataset is currently not allowed.

### v2

@@
 Despite previous efforts, the existing video understanding benchmarks are severely constrained by several issues, especially the insufficient lengths of videos, a lack of diversity in video types and evaluation tasks, and the inappropriateness for evaluating LVU performances.
-To address the above problems, we propose a new benchmark, called MLVU [^1] MLVU Repository: https://github.com/FlagOpen/FlagEmbedding.
+To address the above problems, we propose a new benchmark, called MLVU [^1] MLVU Repository: https://github.com/JUNJIE99/MLVU.
 (Multi-task Long Video Understanding Benchmark), for the comprehensive and in-depth evaluation of LVU.
@@
 1 Introduction
-Large language models (LLMs) are growing into a general solution for numerous AI tasks [6; 39].
+Large language models (LLMs) are growing into a general solution for numerous AI tasks [6; 40].
 In recent years, it becomes increasingly emphasized to extend LLMs with multi-modal capabilities and thus bring the Multi-modal LLM, namely, MLLM.
-Remarkably, it has been made possible for today’s MLLMs to perceive information in texts, images, videos, etc., and solve complicated problems in physical environments [1; 38].
-Along with the development of MLLMs, new benchmarks are continuously created to facilitate comprehensive and in-depth analysis of MLLMs [50; 27; 9; 21].
+Remarkably, it has been made possible for today’s MLLMs to perceive information in texts, images, videos, etc., and solve complicated problems in physical environments [1; 39].
+Along with the development of MLLMs, new benchmarks are continuously created to facilitate comprehensive and in-depth analysis of MLLMs [51; 27; 9; 21].
 However, it remains a great challenge to evaluate the MLLMs’ long-video understanding (LVU) performances given the following limitations.
-Firstly, the majority of existing video understanding benchmarks are made up of short videos [47; 21; 19; 31; 15], whose lengths can be merely a few seconds.
+Firstly, the majority of existing video understanding benchmarks are made up of short videos [48; 21; 19; 32; 15], whose lengths can be merely a few seconds.
 As a result, they are insufficient to reflect the MLLMs’ long-video understanding capabilities.
 Secondly, there is a severe lack of diversity for the existing LVU benchmarks in terms of video genres and evaluation tasks.
-On one hand, the existing benchmarks are usually created based on one type of videos, e.g., movies [35; 22].
-On the other hand, it may only contain one specific type of task in each individual benchmark, for example, the captioning task in [47], the temporal perception task in [49], and the action understanding task in [41; 11].
+On one hand, the existing benchmarks are usually created based on one type of videos, e.g., movies [36; 22].
+On the other hand, it may only contain one specific type of task in each individual benchmark, for example, the captioning task in [48], the temporal perception task in [50], and the action understanding task in [42; 11].
 Due to the above limitations, existing benchmarks have difficulty in performing a comprehensive evaluation of LVU performances.
@@
 Multimodal large language models (MLLMs) have gained wide interests from both academic and industrial sectors.
-Recent studies have made significant progresses in this field by integrating LLM backbones with visual encoders and adapters, and fine-tuning the entire architecture through visual instruction [25; 55; 7].
-Based on the same philosophy, MLLMs have been further developed for video processing using video instruction datasets and specialized video adapters [52; 28; 20; 48; 23; 21].
+Recent studies have made significant progresses in this field by integrating LLM backbones with visual encoders and adapters, and fine-tuning the entire architecture through visual instruction [25; 56; 7].
+Based on the same philosophy, MLLMs have been further developed for video processing using video instruction datasets and specialized video adapters [53; 29; 20; 49; 23; 21].
 However, most existing models are optimized for short videos, typically under one minute, due to the difficulty in establishing sufficient context for longer videos.
@@
 Methods like MovieChat [36] and MA-LMM [12] introduce specialized memory components for recursive video processing.
-Additionally, it is also explored to make selective usage of frames or clips from long videos based on retrievers or agents [46; 33; 40].
+Additionally, it is also explored to make selective usage of frames or clips from long videos based on retrievers or agents [47; 34; 41].
 Despite these progresses, it remains an open problem for MLLMs to effectively handle long videos.
@@
 In video understanding, the research community has made significant efforts as well, particularly for short videos.
-There are specialized benchmarks for temporal perception [49; 42], action understanding [42; 41], video classification [14], video reasoning [45; 44], and video captioning [47; 30].
+There are specialized benchmarks for temporal perception [50; 43; 28], action understanding [43; 42], video classification [14], video reasoning [46; 45], and video captioning [48; 31].
 Recently, MVBench [21] provides a comprehensive short-video benchmark to evaluate general capabilities via question-answering.
@@
 MLVU offers a comprehensive collection of videos of different categories (Figure 1 Middle).
-There are typical real-world videos, which includes movies [51; 35], documentaries [35], TV series, egocentric videos [11], and surveillance footage [37].
+There are typical real-world videos, which includes movies [52; 36], documentaries [36], TV series, egocentric videos [11], and surveillance footage [38].
 Meanwhile, there are also important simulated videos from animated series and game videos [8].
@@
 For example, it only ends up with 64.8% in dealing with the needle question-answering (NQA) task.
-Whereas its analogue tasks in the text domain, e.g., NIHS (Needle-In-the-HayStack-Search) and Passkey Retrieval, can be effectively conquered by many of the existing long LLMs [10; 53].
+Whereas its analogue tasks in the text domain, e.g., NIHS (Needle-In-the-HayStack-Search) and Passkey Retrieval, can be effectively conquered by many of the existing long LLMs [10; 54].
 At the same time, it exhibits even less reliability given tasks like ego-reasoning (ER), action ordering (AO), and action count (AC), while other baseline methods produce even worse performances in these scenarios.
@@
   • K: Datasheet.
-  • Checklist
 Appendix B Collecting Details of our Universal Long Video Collection (ULVC)
@@
 Hosting.
-MLVU can be viewed and downloaded on GitHub at https://github.com/FlagOpen/FlagEmbedding/tree/master/MLVU or on Huggingface at https://huggingface.co/datasets/MLVU/MVLU.
+MLVU can be viewed and downloaded on GitHub at https://github.com/JUNJIE99/MLVU or on Huggingface at https://huggingface.co/datasets/MLVU/MVLU.
 We assure its long-term preservation for future reference and use.
@@
 How will the dataset will be distributed (e.g., tarball on website, API, GitHub)?
-Answer: The benchmark is available on GitHub at https://github.com/FlagOpen/FlagEmbedding/tree/master/MLVU or on Huggingface at https://huggingface.co/datasets/MLVU/MVLU.
+Answer: The benchmark is available on GitHub at https://github.com/JUNJIE99/MLVU or on Huggingface at https://huggingface.co/datasets/MLVU/MVLU.
 Will the dataset be distributed under a copyright or other intellectual property (IP) license, and/or under applicable terms of use (ToU)?
@@
 Will the dataset be updated (e.g., to correct labeling errors, add new instances, delete instances)?
+Answer: Yes. We will post new update in https://github.com/JUNJIE99/MLVU and https://huggingface.co/datasets/MLVU/MVLU if there is any.
 If the dataset relates to people, are there applicable limits on the retention of the data associated with the instances (e.g., were individuals in question told that their data would be retained for a fixed period of time and then deleted)?
@@
 Will older versions of the dataset continue to be supported/hosted/maintained?
+Answer: Yes. Old versions will also be hosted in https://github.com/JUNJIE99/MLVU and https://huggingface.co/datasets/MLVU/MVLU.
 If others want to extend/augment/build on/contribute to the dataset, is there a mechanism for them to do so?

### v3

@@
-MLVU: A Comprehensive Benchmark for Multi-Task Long Video Understanding
+MLVU: Benchmarking Multi-task Long Video Understanding
 Abstract
@@
 Despite previous efforts, the existing video understanding benchmarks are severely constrained by several issues, especially the insufficient lengths of videos, a lack of diversity in video types and evaluation tasks, and the inappropriateness for evaluating LVU performances.
-To address the above problems, we propose a new benchmark, called MLVU [^1] MLVU Repository: https://github.com/JUNJIE99/MLVU.
-(Multi-task Long Video Understanding Benchmark), for the comprehensive and in-depth evaluation of LVU.
+To address the above problems, we propose a new benchmark called MLVU (Multi-task Long Video Understanding Benchmark) for the comprehensive and in-depth evaluation of LVU.
 MLVU presents the following critical values: 1) The substantial and flexible extension of video lengths, which enables the benchmark to evaluate LVU performance across a wide range of durations.
@@
 3) The development of diversified evaluation tasks, which enables a comprehensive examination of MLLMs’ key abilities in long-video understanding.
-The empirical study with 20 latest MLLMs reveals significant room for improvement in today’s technique, as all existing methods struggle with most of the evaluation tasks and exhibit severe performance degradation when handling longer videos.
-Additionally, it suggests that factors such as context length, image-understanding quality, and the choice of LLM backbone can play critical roles in future advancements.
+The empirical study with 23 latest MLLMs reveals significant room for improvement in today’s technique, as all existing methods struggle with most of the evaluation tasks and exhibit severe performance degradation when handling longer videos.
+Additionally, it suggests that factors such as context length, image-understanding ability, and the choice of LLM backbone can play critical roles in future advancements.
 We anticipate that MLVU will advance the research of long video understanding by providing a comprehensive and in-depth analysis of MLLMs.
 1 Introduction
-Large language models (LLMs) are growing into a general solution for numerous AI tasks [6; 40].
+Large language models (LLMs) are growing into a general solution for numerous AI tasks [6, 45].
 In recent years, it becomes increasingly emphasized to extend LLMs with multi-modal capabilities and thus bring the Multi-modal LLM, namely, MLLM.
-Remarkably, it has been made possible for today’s MLLMs to perceive information in texts, images, videos, etc., and solve complicated problems in physical environments [1; 39].
-Along with the development of MLLMs, new benchmarks are continuously created to facilitate comprehensive and in-depth analysis of MLLMs [51; 27; 9; 21].
+Remarkably, it has been made possible for today’s MLLMs to perceive information in texts, images, videos, etc., and solve complicated problems in physical environments [1, 44].
+Along with the development of MLLMs, new benchmarks are continuously created to facilitate comprehensive and in-depth analysis of MLLMs [57, 32, 12, 26].
 However, it remains a great challenge to evaluate the MLLMs’ long-video understanding (LVU) performances given the following limitations.
-Firstly, the majority of existing video understanding benchmarks are made up of short videos [48; 21; 19; 32; 15], whose lengths can be merely a few seconds.
+Firstly, the majority of existing video understanding benchmarks are made up of short videos [52, 26, 22, 36, 19], whose lengths can be merely a few seconds.
 As a result, they are insufficient to reflect the MLLMs’ long-video understanding capabilities.
-Secondly, there is a severe lack of diversity for the existing LVU benchmarks in terms of video genres and evaluation tasks.
-On one hand, the existing benchmarks are usually created based on one type of videos, e.g., movies [36; 22].
-On the other hand, it may only contain one specific type of task in each individual benchmark, for example, the captioning task in [48], the temporal perception task in [50], and the action understanding task in [42; 11].
-Due to the above limitations, existing benchmarks have difficulty in performing a comprehensive evaluation of LVU performances.
+Secondly, there is a notable lack of diversity in both video genres and evaluation tasks.
+Existing benchmarks often concentrate on a single video type, such as egocentric videos [15, 34], or focus on one specific task, like captioning [52].
+These limitations hinder comprehensive evaluation of LVU capabilities.
 Last but not least, many previous evaluation tasks are not properly designed for LVU, as they can be solved without using the complex information from long videos.
-For example, many questions are simply about one single frame in the long videos [36].
-Besides, numerous others are about popular movies and celebrities [22], which can be answered directly by MLLMs based on the textual prompts.
-Conceptually, the MLLMs are expected to handle any type of long video and accomplish any type of related tasks.
+For example, many questions are simply about one single frame in the long videos [41, 60].
+Besides, numerous others are about popular movies and celebrities [27, 13], which can be answered directly by MLLMs based on the textual prompts.
+Conceptually, MLLMs are expected to handle any type of long video and accomplish any related tasks.
 Therefore, the evaluation of LVU should emphasize two important properties: length and diversity.
-Based on such a principle, we propose a novel benchmark called MLVU (Mult-task Long Video Understanding Benchmark), which presents the following critical advantages.
+Furthermore, it is crucial that the evaluation tasks are specifically designed to leverage the complex information inherent in long videos, addressing the shortcomings of previous benchmarks.
+Based on these principles, we propose a novel benchmark called MLVU (Mult-task Long Video Understanding Benchmark), which presents the following critical advantages.
+Benchmarks #Videos #QA Pairs Len.
+(s) Close-Ended Open-Ended Various Genres Multi-Level Multi-Dimension Referring QA NExT-QA [50] 1,000 8,564 39.5 TVQA [21] 15,253 15,253 11.2 MSRVTT-QA [52] 2,900 72,821 15.2 MVBench [26] 3,641 4,000 16.0 Movie101 [58] 101 - 6144 EgoSchema [34] 5,063 5,063 180 MovieChat-1K [41] 130 1,950 500 Video-MME [13] 900 2,700 1024 LongVideoBench [49] 3,763 6,678 473 MLVU 1,730 3,102 930
+Table 1: Comparison of MLVU with existing benchmarks, including the number of videos (#Videos), number of QA pairs (#QA pairs), average video length (Len.), presence of Close-Ended tasks, presence of Open-Ended tasks, inclusion of various video genres (Various Genres), coverage of multiple duration levels (Multi-Level), inclusion of multiple dimensions of LVU tasks (Multi-Dimension), and questions involving local information with clear referring context rather than direct timestamps [41] or well-known narrative elements [17, 27] (Referring QA).
+The first block represents short video understanding benchmarks, and the second block represents long video understanding benchmarks.
+denotes work concurrent with MLVU.
   • It makes a substantial extension for the video length.
 MLVU is created based on long videos of diversified lengths, ranging from 3 minutes to 2 hours.
-The average video length is about 12 minutes, which makes it much longer than most of the existing benchmarks.
+The average video length is about 15 minutes, which makes it much longer than most of the existing benchmarks.
 Additionally, each video is further segmented so that evaluation tasks can be created w.r.t.
@@
 Therefore, it is able to flexibly evaluate the MLLMs’ performance across different video lengths.
-  • It covers a wide variety of video genres.
-On one hand, it contains different types of real-world videos, such as movies, documentaries, surveillance videos, and ego-centric videos.
-On the other hand, it also includes typical simulated videos, like games and cartoons.
-As a result, it is able to comprehensively reflect the MLLMs’ performance across different application scenarios.
+  • It encompasses a wide variety of video genres.
+MLVU includes diverse real-world videos, such as movies, life records, and egocentric videos.
+Additionally, it features typical simulated videos like games and cartoons.
+This diversity allows for a comprehensive assessment of MLLMs’ performance across various application scenarios.
   • It introduces diversified evaluation tasks tailored for LVU.
-There are 9 different tasks in MLVU, which jointly examine a wide range of MLLMs’ key abilities, such as reasoning, captioning, recognition, summarization, etc.
-Both multi-choice and free-form generation tasks are included in MLVU, which reflect the MLLMs’ performances in handling different forms of tasks.
-Finally, some of the tasks are designed to leverage global information from entire videos, while others need to utilize proper local information from specific clips.
-Therefore, it can provide a unified perspective for the degree of completeness and nuance in understanding long videos.
-We extensively investigate 20 popular MLLMs with MLVU, which brings in several critical insights.
+MLVU comprises 9 distinct tasks that collectively assess a wide range of MLLMs’ LVU capabilities.
+On one hand, it includes both multiple-choice and open-ended generation tasks, reflecting the models’ performance in handling different task formats.
+On the other hand, some tasks are designed to leverage global information from entire videos, while others require the use of specific local information from certain clips.
+Moreover, all questions involving local information are annotated with unambiguous context, requiring MLLMs to accurately locate or infer the appropriate clips within long videos.
+Table 1 shows that MLVU provides a more comprehensive evaluation of LVU compared to existing and concurrent benchmarks.
+We extensively investigate 23 popular MLLMs with MLVU, which brings in several critical insights.
 Firstly, long-video understanding remains a technically challenging problem for the existing MLLMs.
-While GPT-4o [^1] https://openai.com/index/hello-gpt-4o/ achieves the leading performance in the experiment, it only attains an average score of 64.6% in multi-choice tasks.
+While GPT-4o [^1] https://openai.com/index/hello-gpt-4o/ achieves the leading performance in the experiment, it only attains an average score of 54.5% in multi-choice tasks.
 All methods struggle with tasks requiring fine-grained information from entire videos, such as action counting, ordering, and summarization.
-Additionally, all methods experience significant performance drops as video lengths increase.
-Secondly, a significant gap persists between open-source and proprietary models.
-Although some of the latest open-source MLLMs can handle very long videos and excel at single-image tasks, they notably lag behind GPT-4o in understanding long videos.
-Finally, the empirical results underscore the influential factors in LVU, such as the extension of context length, the improvement of image understanding ability, and the utilization of strong LLM-backbones.
-In addition to the benchmark’s overall conclusion, the individual tasks enable the fine-grained analysis of MLLMs’ performances in each of the specialized aspects.
+Secondly, recent open-source long video MLLMs have made significant strides in LVU [60, 11, 40].
+These advancements have improved the models’ capability to process extended visual sequences, thereby closing the gap with leading proprietary models in recent months.
+Finally, the empirical results underscore influential factors in LVU, such as the extension of context length, the improvement of image understanding ability, and the utilization of strong LLM-backbones.
+In addition to the benchmark’s overall conclusion, individual tasks enable fine-grained analysis of MLLMs’ performances in each specialized aspects.
 Therefore, we anticipate the benchmark to assist in improving MLLMs’ long-video understanding capabilities by providing insights into their current strengths and weaknesses.
 2 Related Work
-Multimodal Large Langugae Models.
-Multimodal large language models (MLLMs) have gained wide interests from both academic and industrial sectors.
-Recent studies have made significant progresses in this field by integrating LLM backbones with visual encoders and adapters, and fine-tuning the entire architecture through visual instruction [25; 56; 7].
-Based on the same philosophy, MLLMs have been further developed for video processing using video instruction datasets and specialized video adapters [53; 29; 20; 49; 23; 21].
+Figure 1: Statistical Overview of our MLVU benchmark. Left: Video genres included in MLVU; Top Right: Distribution of video duration; Bottom Right: Task types and their counts in MLVU.
+Multimodal Large Language Models.
+Multimodal large language models (MLLMs) have attracted significant interest from both academia and industry.
+Recent advancements in this field have been achieved by integrating LLM backbones with visual encoders and adapters, and fine-tuning the entire architecture through visual instruction tuning [29, 63, 8].
+Based on the same philosophy, MLLMs have been further developed for video processing using video instruction datasets and specialized video adapters [59, 33, 25, 54, 28, 26].
 However, most existing models are optimized for short videos, typically under one minute, due to the difficulty in establishing sufficient context for longer videos.
-To address this challenge, researchers have explored compact representations of videos.
-For example, LLaMa-Vid [22] compresses each video frame into two tokens, allowing the model to handle videos several hours long.
+To address this challenge, researchers have explored compact video representations or extended the context length of MLLMs.
+For instance, LLaMa-Vid [27] compresses each video frame into two tokens, enabling the model to handle videos several hours long.
 Methods like MovieChat [41] and MA-LMM [16] introduce specialized memory components for recursive video processing.
-Additionally, it is also explored to make selective usage of frames or clips from long videos based on retrievers or agents [47; 34; 41].
+Furthermore, approaches such as LWM [30], LongVA [60], and Video-XL [40] are designed to extend the context length of MLLMs, facilitating the processing of longer video inputs.
+Additionally, it is also explored to make selective usage of frames or clips from long videos based on retrievers or agents [53, 38, 46].
 Despite these progresses, it remains an open problem for MLLMs to effectively handle long videos.
@@
 In video understanding, the research community has made significant efforts as well, particularly for short videos.
-There are specialized benchmarks for temporal perception [50; 43; 28], action understanding [43; 42], video classification [14], video reasoning [46; 45], and video captioning [48; 31].
+There are specialized benchmarks for temporal perception [56, 48], action understanding [48, 47], video classification [18], video reasoning [51, 50], and video captioning [52, 35].
 Recently, MVBench [26] provides a comprehensive short-video benchmark to evaluate general capabilities via question-answering.
 For long video understanding, people seek to leverage long-form videos, like movies, to create benchmarks.
-For example, the LVU dataset [44] presents movie understanding tasks, such as predicting release years and identifying character relationships.
-Similarly, LLaMA-Vid [22] created a movie question-answering dataset based on MovieNet [13].
-Despite using long videos, many questions are about common-sense knowledge of the movies.
-As a result, they could be directly answered without using the video’s information.
-In contrast, MovieChat [36] makes use of diversified videos and avoids specific character names or plot details in its questions.
-Considering that many of the questions target on exact time segments, it potentially degrades the tasks to short-video or image understanding problems.
-Beyond movies, there are other task-specific benchmarks like EgoSchema [30], which presents video reasoning tasks using first-person footage from Ego4D [11].
-However, these specialized benchmarks only focus on one single aspect of MLLMs, rather than offering a comprehensive analysis of long video understanding.
-Therefore, it remains essential to develop a comprehensive benchmark with carefully designed tasks to effectively evaluate MLLMs’ capabilities in understanding long videos.
+For example, LLaMA-Vid [27] developed a movie question-answering dataset based on MovieNet [17].
+Despite using long videos, many questions focus on well-known narrative elements, allowing them to be answered without analyzing the video’s content.
+In contrast, MovieChat [41] avoids specific character names or plot details in its questions.
+However, since each question provides a specific timestamp, the tasks can be reduced to short-video or image understanding problems.
+Beyond movies, there are task-specific benchmarks like EgoSchema [34], which presents video reasoning tasks using first-person footage from Ego4D [15].
+These specialized benchmarks, however, focus on a single aspect of MLLMs rather than offering a comprehensive analysis of long video understanding.
+Therefore, it is essential to develop a comprehensive benchmark with carefully designed tasks to effectively evaluate MLLMs’ capabilities in understanding long videos.
 3 MLVU: Multi-task Long Video Understanding Benchmark
@@
 3.1 Overview
-MLVU is a comprehensive benchmark made up of 2593 evaluation tasks [^2] Only consider the tasks for the entire videos.
-Others for fine-grained video segments are not accounted.
-(shown as the right side of Figure 1), which belong to 9 task categories tailored for long-video understanding.
-These categories include 1) Topic Reasoning, 2) Anomaly Recognition, 3) Video Summarization, 4) Needle Question Answering, 5) Ego Reasoning, 6) Plot Question Answering, 7) Sub-Scene Captioning, 8) Action Count, and 9) Action Order.
+MLVU is a multi-task benchmark consisting of 3,102 questions across 9 categories, specifically designed for long video understanding.
+It is divided into a dev set and a test set, containing 2,593 and 509 questions, respectively.
 The benchmark is distinguished by the following features.
-Figure 1: Statistical overview of our LVBench dataset. Left: Distribution of Video Duration; Middle: Distribution of Source Types for Long Videos; Right: Quantification of Each Task Type.
+Diversified Video Categories.
+MLVU offers a comprehensive collection of videos across various categories (Figure 1 Left).
+These include typical real-world videos such as movies, documentaries, TV series, egocentric videos, life records, sports, tutorials, and surveillance footage.
+Additionally, it features significant simulated videos from animated series and game videos.
 Substantial Extension of Video Length.
-MLVU is made up of videos of diversified lengths, spanning from 3 min to more than 2 hours (Figure 1 Left).
-The average length is about 12 min, which is much longer than the existing LVU benchmarks.
+MLVU is made up of videos of diversified lengths, spanning from 3 min to more than 2 hours (Figure 1 Top Right).
 Besides, each video is further partitioned as incremental segments, e.g., the first 3 min, the first 6 min, and the entire video, where tasks are created for each individual segment.
 Thus, the MLLMs can be flexibly evaluated across different video lengths.
-Diversified video categories.
-MLVU offers a comprehensive collection of videos of different categories (Figure 1 Middle).
-There are typical real-world videos, which includes movies [52; 36], documentaries [36], TV series, egocentric videos [11], and surveillance footage [38].
-Meanwhile, there are also important simulated videos from animated series and game videos [8].
 Diversified Evaluation Tasks.
-MLVU also provides a diversified collection of evaluation tasks, which are closely related to the common visual capabilities of MLLMs, such as reasoning, captioning, recognition, perception, and summarization (Figure 1 Right).
+MLVU also provides a diverse array of evaluation tasks, which are closely aligned with the common visual capabilities of MLLMs, such as reasoning, captioning, recognition, perception, and summarization (Figure 1 Bottom Right).
 All the tasks are tailored for LVU.
@@
 Additionally, both multi-choice and free-form generation tasks are included in MLVU, which help to examine MLLMs’ capabilities in handling different task formats.
-Figure 2: Examples of MLVU.
-There are nine tasks designed to evaluate the holistic, single-detail, and multi-detail LVU capabilities of MLLMs.
-The MLLMs are asked to solve the problem (with the ground-truth answers marked in blue) based on the long video input and textual prompt.
 3.2 Construction of MLVU
-The evaluation tasks of MLVU which can be categorized into three types: 1) holistic LVU, which needs to be solved by making use of the global information from the entire video; 2) single-detail LVU, which needs to leverage one critical plot within the long video; 3) multi-detail LVU, which calls for the joint utilization of multiple plots within the long video.
-The construction process of MLVU is discussed w.r.t.
-the above three categories.
-To facilitate the discussion, we define ULVC (Universal Long Video Collection) as the universal collection of miscellaneous long videos, which includes movies, documentaries, TV-series, ego-centric videos, surveillance footages, animated series, game videos, et al.
-(more details about ULVC is presented in the Appendix B).
+The evaluation tasks of MLVU can be categorized into three types: 1) holistic LVU, which needs to be solved by making use of the global information from the entire video; 2) single-detail LVU, which relies on leveraging one critical plot within the long video; and 3) multi-detail LVU, which necessitates the joint utilization of multiple plots within the long video.
+The construction process of MLVU is discussed w.r.t the above three categories.
+To facilitate the discussion, we define ULVC (Universal Long Video Collection) as the universal collection of long videos from various sources (more details about ULVC are presented in Appendix C).
 3.2.1 Holistic LVU
@@
 This includes elements such as the video’s genre, pivotal events, or primary settings.
-Videos for TR tasks are sourced from movies, documentaries, animated series, egocentric videos, and game videos from ULVC.
-All questions and answers undergo manual annotation [^3] Detailed information and annotation guidelines for annotators are presented in Appendix E, resulting in a total of 264 questions.
+All questions and answers undergo manual annotation [^2] Detailed information and annotation guidelines for annotators are presented in Appendix F., resulting in a total of 355 questions.
 TR tasks are formatted as multiple-choice questions, with the model’s performance assessed based on accuracy.
@@
 The selected video clips are longer than three minutes.
-We create 200 questions based on the original annotations provided by the dataset.
+We create 239 questions based on the original annotations provided by the dataset.
 The AR task is also conducted in the multiple-choice format, whose performance is measured by accuracy.
@@
 This task requires MLLMs to summarize the key events in a long video (Figure 2 c).
-We select the narrative-rich videos from ULVC for this task, including movies, TV series, documentaries, and animated series.
-There are 217 selected videos in total, whose summaries are manually annotated.
+We select the narrative-rich videos from ULVC for this task, including movies, TV series, documentaries, life records, and animated series.
+There are 257 selected videos in total, whose summaries are manually annotated.
 During evaluation, the MLLMs are prompted with "Please summarize the main content of this video".
 We employ GPT-4 to assess the generated summaries by comparing with the annotation results.
-Details about annotation and evaluation are presented in Appendix E.3 and F.3.
+Details about annotation and evaluation are presented in Appendix F.3 and G.3.
+Figure 2: Examples of MLVU.
+There are nine tasks designed to evaluate the holistic, single-detail, and multi-detail LVU capabilities of MLLMs.
+The MLLMs are asked to solve the problem (with the ground-truth answers marked in blue) based on the long video input and textual prompt.
+For multiple-choice questions, we set 4 candidates in the dev set and 6 candidates in the test set.
 3.2.2 Single-Detail LVU
@@
 In this task, the MLLM is required to answer a question related to a specific segment (referred as needle) within a long video (referred as background video).
-The needles are short video clips sampled from WebVid [5], while the background videos are sampled from our ULVC.
+The needles are short video clips sampled from WebVid [5] and Clevrer [55], while the background videos are sampled from our ULVC.
 The needle is randomly inserted into the background video, where a question-answer pair is annotated.
@@
 Both videos and QA annotations are collected from the NLQ task of Ego4D [15].
-The ER task is structured as multiple-choice, with a total of 352 questions created for this task.
+The ER task is structured as multiple-choice, with a total of 405 questions created for this task.
 Plot Question-Answering (PQA).
@@
 The video is sampled from the movies, TV series, and animated series in our ULVC.
-There are 539 question-answer pairs created by manual annotation.
+There are 589 question-answer pairs created by manual annotation.
 During annotation, the human annotators are asked to only provide necessary details about the plot but not to suggest any objective hints, e.g., the two characters in the example video are referred as cat and mouse, rather than Tom and Jerry.
-Therefore, it can prevent the question from being short-cut by the MLLM’s common-sense knowledge (more details about PQA can be found in the Appendix E.6).
+Therefore, it can prevent the question from being short-cut by the MLLM’s common-sense knowledge (more details about PQA can be found in the Appendix F.6).
 Sub-Scene Captioning (SSC).
@@
 During evaluation, we employ GPT-4 [1] to measure the quality of caption in comparison with the ground-truth.
-Details about annotation and evaluation are presented in Appendix E.7 and F.3.
+Details about annotation and evaluation are presented in Appendix F.7 and G.3.
 3.2.3 Multi-Detail LVU
@@
 The probes are inserted into a long background video, which is sampled from ULVC.
-There are 259 AO questions in total.
+There are 329 AO questions in total.
 The task is structured as a multiple-choice prblem, where the right order is selected from the misleading options provided by the annotator.
@@
 We also perform manual examination to ensure that the inserted action does not exist in the original background video.
-A total of 206 evaluation instances have been created.
+A total of 266 evaluation instances have been created.
 The AC task is structured as a multiple-choice problem, with performance measured by accuracy.
+Methods Date Input Holistic Single Detail Multi Detail M-Avg G-Avg
+TR AR VS NQA ER PQA SSC AO AC
+Full mark 100 100 10 100 100 100 10 100 100 100 10
+Random 16.7 16.7 16.7 16.7 16.7 16.7 16.7 16.7
+Image MLLMs
+Otter-I [23] 2023-05 16 frm 17.6 17.9 2.03 16.7 17.0 18.0 3.90 15.7 16.7 17.1 2.97
+LLaVA-1.6 [29] 2024-01 16 frm 63.7 17.9 2.00 13.3 26.4 30.0 4.20 21.4 16.7 27.1 3.10
+InternVL-2 [8] 2024-07 16 frm 85.7 51.3 2.55 48.3 47.2 52.0 5.25 32.9 15.0 47.5 3.90
+Claude-3-Opus [2] 2024-03 16 frm 53.8 30.8 2.83 14.0 17.0 20.0 3.67 10.0 6.7 21.8 3.25
+Qwen-VL-Max [4] 2024-01 16 frm 75.8 53.8 3.00 15.0 26.4 4.84 20.0 20.7 11.7 32.2 3.92
+Short Video MLLMs
+Otter-V [23] 2023-05 16 frm 16.5 12.8 2.18 16.7 22.6 22.0 4.20 12.9 13.3 16.7 3.19
+mPLUG-Owl-V [54] 2023-04 16 frm 25.3 15.4 2.20 6.7 13.2 22.0 5.01 14.3 20.0 16.7 3.61
+VideoChat [25] 2023-05 16 frm 26.4 12.8 2.15 18.3 17.0 22.0 4.90 15.7 11.7 17.7 3.53
+Video-LLaMA-2 [59] 2024-08 16 frm 52.7 12.8 2.23 13.3 17.0 12.0 4.87 15.7 8.3 18.8 3.55
+VideoChat2-HD [26] 2024-06 16 frm 74.7 43.6 2.83 35.0 34.0 30.0 5.14 21.4 23.3 37.4 3.99
+Video-LLaVA [28] 2023-11 8 frm 70.3 38.5 20.9 2.30 26.4 26.0 5.06 20.0 21.7 29.3 3.68
+ShareGPT4Video [7] 2024-05 16 frm 73.6 25.6 2.53 31.7 45.3 38.0 4.72 17.1 8.3 34.2 3.63
+VideoLLaMA2 [9] 2024-06 16 frm 80.2 53.8 2.80 36.7 54.7 54.0 5.09 42.9 16.7 48.4 3.95
+Long Video MLLMs
+MovieChat [41] 2023-07 2048 frm 18.7 10.3 2.30 23.3 15.1 16.0 3.24 17.1 15.0 16.5 2.77
+Movie-LLM [42] 2024-03 1 fps 27.5 25.6 2.10 10.0 11.3 16.0 4.93 20.0 21.7 18.9 3.52
+LLaMA-VID [27] 2023-11 1 fps 20.9 23.1 2.70 21.7 11.3 16.0 4.15 18.6 15.0 18.1 3.43
+MA-LMM [16] 2024-04 1000 frm 44.0 23.1 3.04 13.3 30.2 14.0 4.61 18.6 13.3 22.4 3.83
+MiniGPT4-Video [3] 2024-04 90 frm 64.9 46.2 2.50 20.0 30.2 30.0 4.27 15.7 15.0 31.7 3.39
+LongVA [60] 2024-06 256 frm 81.3 41.0 2.90 46.7 39.6 46.0 4.92 17.1 23.3 42.1 3.91
+Video-CCAM [11] 2024-08 96 frm 79.1 38.5 2.65 45.0 52.8 56.0 4.49 24.3 26.7 46.1 3.57
+Video-XL [40] 2024-09 256 frm 78.0 28.2 3.40 50.0 41.5 46.0 5.02 48.6 31.7 46.3 4.21
+LLaVA-Onevision [24] 2024-08 32 frm 83.5 56.4 3.75 46.7 58.4 58.0 5.09 35.7 23.3 51.7 4.42
+GPT-4o [37] 2024-05 0.5 fps 83.7 68.8 4.94 42.9 47.8 57.1 6.80 46.2 35.0 54.5 5.87
+[…]
 4 Experiments and Analysis
 4.1 Settings
-We perform a comprehensive investigation of 20 MLLMs based on MLVU, both open-source and proprietary.
-The experimental MLLMs can be partitioned into three categories.
-1.
-Image MLLMs, which are primarily fine-tuned by image-related instructions.
-We consider the following models for this category: Otter-I [18], LLaVA-1.6 [24], InternVL-1.5 [7], Claude3-Opus [2], Qwen-VL-Max [4], and GPT-4 Turbo [1]).
-2.
-Short Video MLLMs, which are fine-tuned by short-video related instructions.
-This category includes: Otter-V [18], mPlug-Owl-V [49], Video LLaMA-2 [53], Video ChatGPT [29], VideoChat [20], VideoChat2 [21], and Video-LLaVA [23].
-3.
-Long Video MLLMs, which are optimized for their long-video understanding capability.
-This category includes: MiniGPT4-Video [3], LLaMA-VID [22], Movie-LLM [37], MA-LMM [12], MovieChat [36], TimeChat [35], and GPT-4o [33].
+We conduct a comprehensive investigation of 23 MLLMs using our MLVU benchmark, encompassing both open-source and proprietary models.
+The experimental MLLMs are divided into three categories: 1) Image MLLMs, primarily fine-tuned using image-related instructions; 2) Short Video MLLMs, fine-tuned with short-video related instructions; and 3) Long Video MLLMs, optimized for long-video understanding capability.
 For Image MLLMs, we leverage their multi-image inference capabilities to process segmented frames from original videos.
-In the case of Video MLLMs, we employ either a uniform sampling strategy or a frame rate sampling strategy for video processing.
-All models are evaluated based on their official implementations or available APIs, where the evaluation is conducted in a zero-shot manner.
-More details about the evaluation are provided in the Appendix F.
-4.2 Overall Performance
-The overall evaluation results for all investigated MLLMs are shown in Table 1.
-The individual performance is reported for each individual task; meanwhile, the average performances are reported for the multiple-choice (M-Avg) and generation tasks (G-Avg), respectively.
-We can obtain two primary observations from the demonstrated results.
-On one hand, the recently released GPT-4o [33] achieves an overwhelming advantage in our benchmark.
-It outperforms the rest of methods by a big margin in terms of the average performances, with an M-Avg of 64.6% (within 0-100%) and a G-Avg of 5.80 (within 0.0-10.0).
-Besides, it also maintains the leading position in every individual task.
-As for the open-source models, InternVL-1.5 [7] achieves the highest average performance in multiple-choice tasks, whose M-Avg scores reaches 50.4%.
-Such a performance is impressive knowing that it even slightly goes beyond GPT-4 Turbo [1].
-Meanwhile, LLaMA-VID [22] presents the top performance for the generation tasks, whose G-Avg score reaches 4.22.
-However, such a score is significantly lagging behind GPT-4 Turbo and GPT-4o.
-On the other hand, although GPT-4o achieves a huge advantage in comparison with other methods, it actually struggles to handle most of the tasks in the benchmark.
-For example, it only ends up with 64.8% in dealing with the needle question-answering (NQA) task.
-Whereas its analogue tasks in the text domain, e.g., NIHS (Needle-In-the-HayStack-Search) and Passkey Retrieval, can be effectively conquered by many of the existing long LLMs [10; 54].
-At the same time, it exhibits even less reliability given tasks like ego-reasoning (ER), action ordering (AO), and action count (AC), while other baseline methods produce even worse performances in these scenarios.
-The above observations indicate that long-video understanding remains a tough challenge for today’s MLLMs.
+For Video MLLMs, we employ either a uniform sampling strategy or a frame rate sampling strategy for video processing.
+All models are evaluated based on their official implementations or available APIs, with evaluations conducted in a zero-shot manner.
+More details about the evaluation are provided in Appendix G.
+4.2 Main Results
+The overall evaluation results for all investigated MLLMs in the MLVU test set are shown in Table 2 (with dev set results in Appendix B).
+Individual performances are reported for each task, while average performances are provided for multiple-choice (M-Avg) and generation tasks (G-Avg).
+From the results, we derive three primary conclusions:
+1) The proprietary model GPT-4o [37] achieves optimal performance in our benchmark.
+It leads in multiple-choice tasks with an M-Avg of 54.5%(within 0-100%) and excels in generation tasks with a G-Avg of 5.87 (within 0.0-10.0), outperforming all other methods.
+2) Recent advances in LVU have achieved significant progress, and the gap between open-source long video MLLMs and GPT-4o on close-ended tasks is narrowing.
+Before June 2024, the best open-source long video MLLMs, MiniGPT4-Video [3], lagged significantly behind GPT-4o.
+However, recent models [60, 11, 40, 24] have made substantial progress.
+For instance, LLaVA-Onevision trails GPT-4o by only 2.8% in M-Avg.
+These models have improved their ability to handle long visual sequences, achieving significant advancements in single-detail (e.g., NQA) and multi-detail (e.g., AC) tasks compared to previous open-source models.
+3) Existing methods still struggle to handle most tasks in our benchmark.
+For instance, GPT-4o only achieves 42.9% in the needle question-answering (NQA) task.
+In contrast, analogous tasks in the text domain, such as NIHS (Needle-In-the-Haystack-Search) and Passkey Retrieval, are effectively handled by many existing long LLMs [14, 61].
+Additionally, GPT-4o shows even less reliability in tasks like ego-reasoning (ER), action ordering (AO), and action count (AC), with most baseline methods performing even worse.
+These observations indicate that long-video understanding remains a significant challenge for today’s MLLMs.
 In addition to the primary conclusions from the overall performances, we can also make the following interesting observations about the individual tasks.
-First of all, the multiple-choice holistic tasks, i.e., topic retrieval (TR) and anomaly recognition (AR), present much higher differentiation than other tasks.
-Proprietary MLLMs, like GPT-4o, GPT-4-turbo, and superior open-source models, like InternVL-1.5, can accurately solve such problems; meanwhile, many other popular MLLMs still fail to generate meaningful performances.
-Knowing that the two tasks only require an overall understanding of the long videos, they can serve as a preliminary indicator of MLLMs’ LVU ability.
-Besides, it’s hard to deal with tasks which need nuanced understanding of multiple details.
+4) The close-ended holistic tasks present much higher differentiation than other tasks.
+These tasks, i.e., topic reasoning (TR) and anomaly recognition (AR), show significant variance in performance across different models.
+Proprietary MLLMs, like GPT-4o, and superior open-source models, such as InternVL-2 [8], VideoLLaMA2 [9], and LLaVA-OneVision [24], can accurately solve these problems.
+Meanwhile, many other popular MLLMs still fail to generate meaningful performances.
+Since these tasks only require an overall understanding of long videos, they can serve as a preliminary indicator of MLLMs’ long video understanding (LVU) ability.
+5) It’s challenging to deal with tasks that require nuanced understanding of multiple details.
 Although several MLLMs can handle single-detail LVU tasks to some extent, their performances suffer from catastrophic degradation when addressing multi-detail LVU tasks.
-Most methods, except for GPT-4o, fail entirely in action order (AO) and action count (AC) tasks.
-Additionally, most approaches struggle with the summarization task, which require recalling multiple nuanced details from long videos.
-Finally, the existing long-videos MLLMs (except for GPT-4o) are no better than those primarily trained from short-video or image related instructions.
-Although such models can intake much longer videos (i.e.
-much more frames than other baselines), it turns out that the extended input information is not effectively utilized.
-On the contrary, Video-LLaVA can achieve a relatively competitive performance, especially in multiple-choice, despite that it can only intake 8 frames.
+Most methods, except for GPT-4o and Video-XL [40], fail entirely in action order (AO) and action count (AC) tasks.
+Additionally, most approaches struggle with summarization tasks, which require recalling multiple nuanced details from long videos.
 As a brief conclusion, although today’s MLLMs can deal with some preliminary LVU tasks, it remains a tough challenge to achieve an in-depth understanding of nuanced information within long videos.
-Methods Input Holistic Single Detail Multi Detail M-Avg G-Avg
-TR AR VS NQA ER PQA SSC AO AC
-Full mark - 100 100 10 100 100 100 100 10 100 100 10
-Image MLLMs
-[Uncaptioned image] Otter-I [18] 16 frm 25.0 25.0 2.18 25.1 25.0 24.9 4.12 13.1 25.2 23.3 3.15
-[Uncaptioned image]LLaVA-1.6 [25] 16 frm 60.6 41.0 2.11 43.1 38.4 41.0 4.35 25.5 25.7 39.3 3.23
-[Uncaptioned image]InternVL-1.5 [7] 16 frm 78.8 67.0 3.16 52.7 43.5 54.4 4.88 32.8 23.8 50.4 4.02
-[Uncaptioned image]Claude-3-Opus [2] 16 frm 67.2 43.5 3.11 21.6 40.2 47.8 3.66 18.2 16.7 36.5 3.39
-[Uncaptioned image]Qwen-VL-Max [4] 16 frm 67.4 63.5 2.71 40.3 40.9 43.3 5.21 25.0 14.8 42.2 3.96
-[Uncaptioned image] GPT-4 Turbo [1] 16 frm 79.5 68.0 4.49 45.9 47.4 60.6 6.21 26.5 16.1 49.2 5.35
-Short Video MLLMs
-[Uncaptioned image]Otter-V [18] 16 frm 24.6 26.0 2.38 28.2 27.6 22.3 4.23 15.1 26.7 24.4 3.31
-[Uncaptioned image]mPLUG-Owl-V [49] 16 frm 28.0 25.0 2.36 24.5 31.8 27.3 5.31 21.2 23.3 25.9 3.84
-[Uncaptioned image]VideoChat [20] 16 frm 33.0 32.0 2.31 27.0 32.1 27.6 5.01 24.3 28.6 29.2 3.66
-[Uncaptioned image]Video-ChatGPT [29] 100 frm 26.9 24.0 2.31 40.3 42.0 29.9 5.48 25.1 31.1 31.3 3.90
-[Uncaptioned image]Video-LLaMA-2 [53] 16 frm 54.5 41.5 2.34 39.4 33.5 35.4 5.22 18.5 25.7 35.5 3.78
-[Uncaptioned image]VideoChat2 [21] 16 frm 74.6 51.5 2.57 42.0 47.4 43.8 5.04 22.8 29.6 44.5 3.81
-[Uncaptioned image]Video-LLaVA [23] 8 frm 71.6 57.0 2.43 53.2 45.2 48.4 5.25 20.1 35.9 47.3 3.84
-Long Video MLLMs
-[Uncaptioned image]MovieChat [36] 2048 frm 29.5 25.0 2.33 24.2 24.7 25.8 3.23 28.6 22.8 25.8 2.78
-[Uncaptioned image]Movie-LLM [37] 1 fps 30.0 29.0 2.88 29.6 24.7 24.1 5.00 20.5 24.8 26.1 3.94
-[Uncaptioned image]TimeChat [35] 96 frm 23.1 27.0 2.54 24.5 28.4 25.8 4.29 24.7 32.0 30.9 3.42
-[Uncaptioned image]LLaMA-VID [22] 1 fps 50.8 34.5 3.22 30.1 32.7 32.5 5.22 23.9 27.8 33.2 4.22
-[Uncaptioned image]MA-LMM [12] 1000 frm 51.9 35.5 2.12 43.1 38.9 35.8 4.80 25.1 24.3 36.4 3.46
-[Uncaptioned image] MiniGPT4-Video [3] 90 frm 70.9 52.5 2.64 49.0 48.6 44.5 4.07 23.2 23.0 44.5 3.36
-[Uncaptioned image]GPT-4o [33] 0.5 fps 87.4 74.5 4.90 64.8 57.1 65.1 6.69 56.7 46.3 64.6 5.80
-[…]
-Figure 3: Experimental performances on MLVU under various video lengths: GPT4-turbo (orange), MiniGPT4-Video (red), VideoChat2 (green), Video-LLaMA-2 (blue), and MA-LMM (cyan).
-4.3 Detailed Analysis
-We analyze the impact from video length and three factors: context length, image understanding (IU) ability, LLM-backbone. All the factors are empirically critical to MLLMs’ LVU performances.
-In the first place, we evaluate MLLMs’ performances across various video lengths.
+4.3 Further Analysis
+     Impact of Context Length Impact of IU Impact of LLM
+  Model Context Len. M-Avg Model MMMU (Val) M-Avg Model LLM M-Avg
+  MGV 16 24.2 Otter-I 32.2 17.1 VLM2 Vicuna-7B 13.3
+90 31.7$\uparrow$7.5 LLaVA-1.6 35.8 27.1$\uparrow$10.0 Vicuna-13B 18.8$\uparrow$5.5
+GPT-4o 16 45.8 GPT-4V 58.1 43.3 MGV LLaMA-7B 20.6
+256 54.5$\uparrow$8.7 GPT-4o 63.8 45.8$\uparrow$2.5 Mistral-7B 31.7 $\uparrow$11.1
+Table 3: Detailed discussions about the impact from context length, image understanding (IU) ability, and LLM Backbone.
+For the IU impact experiment, we used 16-frame uniform sampling for both GPT-4V and GPT-4o.
+MGV: MiniGPT4-Video, VLM2: Video-LLaMA-2.
+6) Longer videos are more challenging for MLLMs.
+We evaluate MLLMs’ performances across various video lengths.
 For this purpose, we introduce a derivative dataset alongside MLVU, called MLVU Time-ladder.
@@
 As shown in Figure 3, the performances of all models tend to decline as the video length grows, which indicates that the existing MLLMs’ LVU abilities are severely constrained by the video length.
-By comparison, image models, like GPT-4-turbo [1], and short-video models, like VideoChat2 [21], VideoLLaMA [53], are more vulnerable to the growth of video length, while long video models, like MiniGPT4-Video, can be relatively more resilient.
-     Impact of Context Length Impact of IU Impact of LLM
-  Model Context Len. M-Avg Model MMMU (Val) M-Avg Model LLM M-Avg
-  MGV 16 41.1 Otter 32.2 24.4 VLM2 Vicuna-7B 24.3
-90 44.5$\uparrow$3.4 LLaVA 35.8 39.3$\uparrow$14.9 Vicuna-13B 35.5$\uparrow$11.2
-GPT-4o 16 54.2 GPT4V 58.1 49.2 MGV LLAMA-7B 33.7
-256 64.6$\uparrow$10.4 GPT-4o 63.8 54.2$\uparrow$5.0 Mistral-7B 44.5 $\uparrow$10.8
-Table 2: Detailed analysis about the impact from context length, image understanding (IU) ability, and LLM Backbone. MGV: MiniGPT4-Video, GPT4V: GPT-4-turbo, VLM2: Video-LLaMA-2.
-We further examine MLLMs’ performances (measured by M-Avg) under varying context lengths.
-Specifically, we increase MiniGPT4-Video’s input from 16 to 90 frames and GPT-4o’s input from 16 to 256 frames (the left side of Table 2).
-As the input length extends, both models consistently show improved performance.
-To investigate the impact of MLLMs’ image understanding (IU) ability, we reference the experiment results from MMMU [51] (the middle of Table 2).
-It’s evident that the MLLMs’ LVU performances basically align with their image understanding performances in MMMU.
-Finally, we compare MLLMs with different backbones (the right side of Table 2).
-The results show that LVU performances improve with larger (Vicuna-13B vs.
-Vicuna-7B) and more capable (Mistral-7B vs.
-Llama-2-7B) backbone encoders.
+Moreover, the short video model Video-LLaMA-2 [59] maintains a certain level of LVU ability at 3 minutes, but its performance approaches random results at 10 minutes.
+7) The performance of recent advanced long video MLLMs remains robust regardless of the position of the referring clip within the long video.
+In single-detail tasks, the referring clip denotes the specific segment of the long video that is referenced or inferred to answer a question.
+As shown in Figure 4, we categorize clip positions into four intervals and assess model performance on two single-detail tasks: ego reasoning (ER) and plot question-answering (PQA).
+Recent long video MLLMs, such as LongVA [60] and Video-XL [40], maintain consistent performance regardless of the referring clip’s position within the video.
+Conversely, short video MLLMs are more sensitive to clip location.
+This indicates that recent advancements in long video MLLMs enhance both reliable clue retrieval and effective reasoning from extended visual sequences.
+Figure 3: Experimental performance on varying video lengths. The evaluated metric is the average accuracy across five multiple-choice tasks involving local information: NQA, ER, PQA, AC, and AO.
+Figure 4: Model performance across different referring clip positions, spanning from the beginning to the end of the entire video.
+8) The challenge of multi-detail tasks increases with the number of details.
+We analyzed model performance on the action count (AC) task by grouping questions based on the number of probes (which correspond to details) and evaluating the average performance within these groups.
+As shown in Figure 5, performance significantly declines across all models as the number of probes increases.
+This indicates that current MLLMs face substantial difficulties comprehending and processing multiple details simultaneously, highlighting a critical area for future improvement in long video understanding capabilities.
+Figure 5: Model performance on the action count (AC) task in relation to the number of probes.
+9) Context Length, Image-Understanding ability, and the choice of LLM Backbones are key factors in LVU performance.
+As shown in Table 3, we conducted ablation experiments on several factors affecting MLLMs, using M-Avg as the evaluation metric.
+First, we examined the models’ handling of different context lengths.
+Specifically, we increased MiniGPT4-Video’s input from 16 to 90 frames and GPT-4o’s input from 16 to 256 frames (as shown on the left side of Table 3).
+Both models showed consistent performance improvements with longer input lengths.
+To assess the impact of image understanding (IU) capabilities, we referred to the results from MMMU [57] (presented in the middle of Table 3).
+It is evident that MLLMs’ LVU performance generally aligns with their IU performance in MMMU.
+Finally, we compared MLLMs using different backbones (depicted on the right side of Table 3).
+The findings indicate that LVU performance improves with larger (Vicuna-13B vs.
+Vicuna-7B) and more advanced backbones (Mistral-7B vs.
+Llama-2-7B).
 These observations indicate that LVU is the result of multiple complex factors, with the ability to perceive longer videos and effectively utilize the perceived information being crucial for the improvement of LVU.
-5 Conclusion & Discussion
+5 Conclusion
 This paper presents MLVU, a novel benchmark for the assessment of long video understanding.
@@
 We anticipate this benchmark will facilitate future research in long-video understanding of MLLMs.
-Limitations.
-While MLVU basically covers the major dimensions of long-video understanding, it can still be improved for even better comprehensiveness.
-For instance, there can be tasks related to high-resolution videos, or more specific tasks such as tracking and low-level processing.
-Our work will be a persistent effort, where new tasks and video types will be continually introduced.
+Supplementary Material
 Appendix A Overview of Appendix
-  • B: Collecting Details of our Universal Long Video Collection (ULVC).
-  • C: Distribution of Video Durations for Each Task.
+  • B: Evaluation Results on MLVU Dev Set
+  • C: Collecting Details of our Universal Long Video Collection (ULVC)
   • D: Details of the MLVU Time-Ladder
-  • E: Annotation Details of MLVU.
-  • F: Details of Baselines and the Evaluation Process.
-  • G: Explorations of Video Retrieval Augmented Generation.
-  • H: More Experimental Results.
-  • I: More Visualized Examples of MLVU.
-  • J: Licensing, Hosting and Maintenance Plan.
-  • K: Datasheet.
-Appendix B Collecting Details of our Universal Long Video Collection (ULVC)
+  • E Detailed Division of Dev and Test Sets in MLVU
+  • F: Annotation Details of MLVU
+  • G: Details of Baselines and the Evaluation Process.
+  • H: Explorations of Video Retrieval Augmented Generation
+  • I: More Visualized Examples of MLVU
+Appendix B Evaluation Results on MLVU Dev Set
+The evaluation results of the baselines on the MLVU dev set are detailed in Table 2.
+Notably, the multiple-choice questions in the MLVU dev set present four options, whereas the MLVU test set offers six, making the latter more challenging and discriminative.
+Appendix C Collecting Details of our Universal Long Video Collection (ULVC)
 In the initial stage of our Multi-task Long Video Understanding (MLVU) benchmark creation, we first collected long-form videos from a variety of sources to form our Universal Long Video Collection (ULVC).
 The entirety of the long videos incorporated into our MLVU benchmark were selected, edited, or synthesized from ULVC.
-Specifically, our ULVC consists of a diverse set of 757 long videos.
-This collection includes 168 movies sourced from the Movie101 dataset [52] and the MovieChat dataset [36], as well as 60 documentaries from the MovieChat dataset [36].
-Additionally, it comprises 65 game videos from the MineDojo dataset [8], 200 surveillance videos from the UCF-Crime dataset [38], and 100 ego-centric videos from the Ego4D dataset [11].
-Furthermore, our team independently collected 72 cartoons and 92 TV series to enrich the dataset.
-It’s important to clarify that the quantity of videos in the ULVC does not directly correspond to the number of videos and questions in our MLVU benchmark, which are 1334 and 2593 respectively.
-For example, a two-hour movie from the ULVC might be utilized in its entirety for the Sub-Scene Captioning task, or it could be segmented into several approximately 10-minute clips for the Video Summary task, or even used as a background video for synthetic video generation.
+Specifically, our ULVC includes a diverse set of 986 long videos.
+This collection features 168 movies from the Movie101 [58] and MovieChat [41] datasets, along with 60 documentaries from MovieChat [41].
+It also contains 65 game videos from MineDojo [10], 239 surveillance videos from UCF-Crime [43], and 100 ego-centric videos from Ego4D [15].
+Additionally, we independently collected 72 cartoons, 92 TV series, 60 tutorial videos, 60 sports videos, and 70 life records.
+It’s important to clarify that the quantity of videos in the ULVC does not directly correspond to the number of videos and questions in our MLVU benchmark, which are 1,730 and 3,102 respectively.
+For example, a two-hour movie from the ULVC might be utilized in its entirety for the Sub-Scene Captioning task, or it could be segmented into several approximately 10-minute clips for the Video Summarization task, or even used as a background video for synthetic video generation.
 Moreover, a single video could be annotated with multiple questions simultaneously.
-Appendix C Distribution of Video Durations for Each Task
-Figure 4 illustrates the detailed video duration distribution for each task in our MLVU, featuring videos of varying lengths from three minutes to over 120 minutes.
-Figure 4: Video duration distributions of different tasks in our MLVU.
 Appendix D Details of the MLVU Time-Ladder
@@
 Lastly, for the synthetic tasks of NQA, AO, and AC, we possess the capability to directly generate the necessary video lengths.
-Appendix E Annotation Details of MLVU
-E.1 Topic Reasoning (TR).
+Appendix E Detailed Division of Dev and Test Sets in MLVU
+Our MLVU comprises a total of 3,102 questions, divided into a dev set with 2,593 questions and a test set with 509 questions.
+We present the detailed distribution of questions for each task in Table 1.
+Task Dev Test Total
+Topic Reasoning 264 91 355
+Anomaly Recognition 200 39 239
+Video Summarization 217 40 257
+Needle QA 355 60 415
+Ego Reasoning 352 53 405
+Plot QA 539 50 589
+Sub-Scene Captioning 201 46 247
+Action Order 259 70 329
+Action Count 206 60 266
+Table 1: Detailed Distribution of Questions in the MLVU Dataset Across Dev and Test Sets for Each Task.
+Methods Date Input Holistic Single Detail Multi Detail M-Avg G-Avg
+TR AR VS NQA ER PQA SSC AO AC
+Full mark 100 100 10 100 100 100 10 100 100 100 10
+Random 25 25 25 25 25 25 25 25
+Image MLLMs
+Otter-I [23] 2023-05 16 frm 25.0 25.0 2.18 25.1 25.0 24.9 4.12 13.1 25.2 23.3 3.15
+LLaVA-1.6 [29] 2024-01 16 frm 60.6 41.0 2.11 43.1 38.4 41.0 4.35 25.5 25.7 39.3 3.23
+Claude-3-Opus [2] 2024-03 16 frm 67.2 43.5 3.11 21.6 40.2 47.8 3.66 18.2 16.7 36.5 3.39
+Qwen-VL-Max [4] 2024-01 16 frm 67.4 63.5 2.71 40.3 40.9 43.3 5.21 25.0 14.8 42.2 3.96
+Short Video MLLMs
+Otter-V [23] 2023-05 16 frm 24.6 26.0 2.38 28.2 27.6 22.3 4.23 15.1 26.7 24.4 3.31
+mPLUG-Owl-V [54] 2023-04 16 frm 28.0 25.0 2.36 24.5 31.8 27.3 5.31 21.2 23.3 25.9 3.84
+VideoChat [25] 2023-05 16 frm 33.0 32.0 2.31 27.0 32.1 27.6 5.01 24.3 28.6 29.2 3.66
+Video-LLaMA-2 [59] 2024-08 16 frm 54.5 41.5 2.34 39.4 33.5 35.4 5.22 18.5 25.7 35.5 3.78
+VideoChat2-HD [26] 2024-06 16 frm 74.6 51.5 2.57 42.0 47.4 43.8 5.04 22.8 29.6 44.5 3.81
+Video-LLaVA [28] 2023-11 8 frm 71.6 57.0 2.43 53.2 45.2 48.4 5.25 20.1 35.9 47.3 3.84
+ShareGPT4Video [7] 2024-05 16 frm 75.8 51.5 2.52 47.6 43.2 48.4 5.02 34.0 23.3 46.4 3.77
+VideoLLaMA2 [9] 2024-06 16 frm 74.6 64.5 2.79 49.9 43.8 45.1 5.18 34.0 27.4 48.5 3.99
+Long Video MLLMs
+MovieChat [41] 2023-07 2048 frm 29.5 25.0 2.33 24.2 24.7 25.8 3.23 28.6 22.8 25.8 2.78
+Movie-LLM [42] 2024-03 1 fps 30.0 29.0 2.88 29.6 24.7 24.1 5.00 20.5 24.8 26.1 3.94
+TimeChat [39] 2023-12 96 frm 23.1 27.0 2.54 24.5 28.4 25.8 4.29 24.7 32.0 30.9 3.42
+LLaMA-VID [27] 2023-11 1 fps 50.8 34.5 3.22 30.1 32.7 32.5 5.22 23.9 27.8 33.2 4.22
+MA-LMM [16] 2024-04 1000 frm 51.9 35.5 2.12 43.1 38.9 35.8 4.80 25.1 24.3 36.4 3.46
+MiniGPT4-Video [3] 2024-04 90 frm 70.9 52.5 2.64 49.0 48.6 44.5 4.07 23.2 23.0 44.5 3.36
+LongVA [60] 2024-06 256 frm 83.3 58.5 3.39 69.3 50.0 67.2 5.26 38.6 27.2 56.3 4.33
+Video-CCAM [11] 2024-08 96 frm 84.9 66.0 2.84 73.2 60.5 66.1 5.19 42.1 38.4 63.1 4.01
+Video-XL [40] 2024-09 256 frm 80.3 54.5 3.25 73.8 57.4 67.9 5.02 68.3 40.3 64.9 4.14
+GPT-4o [37] 2024-05 0.5 fps 87.4 74.5 4.90 64.8 57.1 65.1 6.69 56.7 46.3 64.6 5.80
+[…]
+Appendix F Annotation Details of MLVU
+F.1 Topic Reasoning (TR).
 The questions and corresponding answers for the TR task were meticulously annotated by human annotators, following the specific guidelines illustrated in Figure 1.
@@
 Figure 1: Annotation Guidelines for the Topic Reasoning Task.
-E.2 Anomaly Recognition (AR).
+F.2 Anomaly Recognition (AR).
 The anomaly recognition task did not involve manual annotation.
@@
 We also modified the original labels to fit a multiple-choice format.
-E.3 Video Summarization (VS).
+F.3 Video Summarization (VS).
 The ground truth data for the VS task were derived from manual annotations.
@@
 Figure 2: Annotation Guidelines for the Video Summarization Task.
-E.4 Needle Question-Answering (NQA).
+F.4 Needle Question-Answering (NQA).
 We leveraged the GPT-4 [1] and the detailed video caption data from the WebVid dataset [5] to facilitate a semi-automated generation of annotated questions and answers for the NQA task.
@@
 Figure 3: The prompt provided to GPT-4 in the process of creating the question-answer pair for the Needle Question-Answering task.
-E.5 Ego Reasoning (ER).
+F.5 Ego Reasoning (ER).
 The video resources, questions, and correct responses used in the ER task were derived from the Natural Language Queries (NLQ) task within the Ego4D dataset [15].
 This data was restructured to fit a multiple-choice question format.
-E.6 Plot Question-Answering (PQA).
+F.6 Plot Question-Answering (PQA).
 The PQA task’s questions and answers were annotated by human annotators, following specific guidelines illustrated in Figure 4.
@@
 Our PQA addresses these issues, providing enhanced usability and reliability.
-E.7 Sub-Scene Captioning (SSC).
+F.7 Sub-Scene Captioning (SSC).
 In the development process of the SSC task, we employed human annotators to generate both prompts and standard caption data.
@@
 To prevent any leakage of information, the prompt was designed to exclude any character-specific names or objective hints, instead incorporating rich descriptive details to allude to the plot.
-Following this, the annotators produced a detailed caption for this sub-scene, and deconstructed the caption into multiple, non-redundant "scoring points" to facilitate quantitative assessment (the details of the evaluation metric can be found in Section F.3).
+Following this, the annotators produced a detailed caption for this sub-scene, and deconstructed the caption into multiple, non-redundant "scoring points" to facilitate quantitative assessment (the details of the evaluation metric can be found in Section G.3).
 More visualized examples of PQA task can be found in Figure 12.
 Figure 6: Annotation Guidelines for the Sub-Scene Captioning Task.
-E.8 Action Order (AO).
+F.8 Action Order (AO).
 The videos, questions, and answers for the action order task were all synthetically generated.
@@
 This rigorous methodology ensured the reliability of our data.
-E.9 Action Count (AC).
+F.9 Action Count (AC).
 The process of data acquisition and annotation for the action count task closely mirrored that of the action order task.
@@
 We employed a strategy consistent with the action order task to ensure the validity and reliability of our evaluation data.
-Appendix F Details of Baselines and the Evaluation Process
-F.1 Baselines
-In this section, we detail the primary baselines evaluated on our MLVU.
-For image-based MLLMs, most available models lack multi-image inference capabilities.
-Consequently, we select Otter-I, LLaVA-1.6, and InternVL, which offer official multi-image implementations.
-Additionally, we include three proprietary models—Claude-3-Opus, Qwen-VL-Max, and GPT-4 Turbo—that provide APIs for multi-image inference.
-For available models, we estimate the maximum input frames based on their maximum LLM context length.
-For Claude and Qwen, the supported maximum image numbers are approximately 20; therefore, we select 16 frames to ensure fair comparisons.
-Similarly, we choose 16 frames for GPT-4 Turbo to maintain consistency.
-In terms of video MLLMs, we adhere to the default settings for frame sampling strategies.
-For instance, VideoChat2 uniformly samples 16 frames, whereas LLaMA-Vid samples 1 frame per second.
-Specifically, the GPT-4o can support a maximum of approximately 500 images when each image’s resolution is set to 512$\times$512 pixels.
-Consequently, we opt for a sampling rate of 0.5 fps to accommodate the majority of our videos.
-Table 3 and Table 4 show the LLM-backbone as well as evaluated weight links used in each MLLM.
-     Image MLLMs Short Video MLLMs Long Video MLLMs
-  Model LLM Model LLM Model LLM
-  Otter-I MPT-7B Otter-V LLaMA-7B Movie-LLM Vicuna-7B
-LLaVA-1.6 Vicuna-7B mPLUG-Owl-V LLaMA-7B MovieChat Vicuna-7B
-InternVL-1.5 Vicuna-13B VideoChat Vicuna-7B TimeChat LLaMA-2 7B
-Claude-3-Opus Claude3 Video-ChatGPT Vicuna-7B LLaMA-VID Vicuna-7B
-Qwen-VL-Max Qwen Video-LLaMA-2 Vicuna-13B MA-LMM Vicuna-7B
-GPT-4 Turbo GPT-4 VideoChat2 Vicuna-7B MiniGPT4-Video Mistral-7B
-Video-LLaVA Vicuna-7B GPT-4o GPT4
-Table 3: LLM-Backbone in different MLLMs
-     Image MLLMs Short Video MLLMs Long Video MLLMs
-  Model Weight Model Weight Model Weight
-  Otter-I Link Otter-V Link Movie-LLM Link
-LLaVA-1.6 Link mPLUG-Owl-V Link MovieChat Link
-InternVL-1.5 Link VideoChat Link TimeChat Link
-Claude-3-Opus - Video-ChatGPT Link LLaMA-VID Link
-Qwen-VL-Max - Video-LLaMA-2 Link MA-LMM Link
-GPT-4 Turbo - VideoChat2 Link MiniGPT4-Video Link
-Video-LLaVA Link GPT-4o -
-Table 4: Link of evaluation weights. The specific version for “GPT-4 Turbo” and “GPT-4o” is “gpt-4-turbo-2024-04-09” and “gpt-4o-2024-05-13”.
-F.2 Inference Detatils
+Appendix G Details of Baselines and the Evaluation Process
+G.1 Baselines
+In this section, we outline the primary baselines evaluated on our MLVU.
+For image-based MLLMs, most models lack multi-image inference capabilities.
+Therefore, we select Otter-I, LLaVA-1.6, and InternVL, which have official multi-image implementations.
+Additionally, we include two proprietary models—Claude-3-Opus and Qwen-VL-Max—that offer APIs for multi-image inference.
+For the available models, we determine the maximum input frames based on their LLM context length.
+Claude and Qwen support a maximum of approximately 20 images, so we choose 16 frames to ensure fair comparisons.
+Regarding video MLLMs, we use default frame sampling strategies.
+For example, VideoChat2 uniformly samples 16 frames, while LLaMA-Vid samples 1 frame per second.
+Specifically, GPT-4o can handle up to approximately 500 images at a resolution of 512×512 pixels.
+Thus, we select a sampling rate of 0.5 fps to accommodate most of our videos.
+G.2 Inference Detatils
 We have developed two templates specifically for Multiple-Choice and Generation tasks, as illustrated in Figure 7.
@@
 In our evaluation, the templates are seamlessly integrated into the evaluation code of open-release models or available API of proprietary models.
-F.3 Evaluation Metrics
+G.3 Evaluation Metrics
 For the evaluation of Multiple Choice tasks, we directly compute absolute accuracy by matching the predicted option with the ground truth.
@@
 Figure 8: Detailed prompt for evaluation of generation tasks in MLVU.
-Appendix G Explorations of Video Retrieval Augmented Generation
+Appendix H Explorations of Video Retrieval Augmented Generation
 As discussed in Section 4.3, most MLLMs are adversely affected by video length.
@@
 We evaluate two image MLLMs and three video MLLMs in different settings.
-Rank Model Score
-  Refer to caption GPT-4o 87.4
-Refer to caption GPT-4 Turbo 79.5
-Refer to caption InternVL-1.5 78.8
-4 VideoChat2 74.6
-5 Video-LLaVA 71.6
-6 MiniGPT4-Video 70.9
-7 Qwen-VL-Max 67.4
-8 Claude-3-Opus 67.2
-9 LLaVA 60.6
-10 Video-LLaMA-2 54.5
-(a) Anomaly Recognition
-Rank Model Score
-  Refer to caption GPT-4o 74.5
-Refer to caption GPT-4 Turbo 68.0
-Refer to caption InternVL-1.5 67.0
-4 Qwen-VL-Max 63.5
-5 Video-LLaVA 57.0
-6 MiniGPT4-Video 52.5
-7 VideoChat2 51.5
-8 Claude-3-Opus 43.5
-9 Video-LLaMA-2 41.5
-10 LLaVA 41.0
-(b) Topic Reasoning
-Rank Model Score
-  Refer to caption GPT-4o 4.90
-Refer to caption GPT-4 Turbo 4.49
-Refer to caption LLaMA-VID 3.22
-4 InternVL-1.5 3.16
-5 Claude-3-Opus 3.11
-6 Movie-LLM 2.88
-7 Qwen-VL-Max 2.71
-8 MiniGPT4-Video 2.64
-9 VideoChat2 2.57
-10 TimeChat 2.54
-(c) Video Summarization
-Rank Model Score
-  Refer to caption GPT-4o 64.8
-Refer to caption Video-LLaVA 53.2
-Refer to caption InternVL-1.5 52.7
-4 MiniGPT4-Video 49.0
-5 GPT-4 Turbo 45.9
-6 LLaVA 43.1
-7 MA-LMM 43.1
-8 VideoChat2 42.0
-9 Qwen-VL-Max 40.3
-10 Video-ChatGPT 40.3
-(d) Needle QA
-Rank Model Score
-  Refer to caption GPT-4o 57.1
-Refer to caption MiniGPT4-Video 48.6
-Refer to caption GPT-4 Turbo 47.4
-4 VideoChat2 47.4
-5 Video-LLaVA 45.2
-6 InternVL-1.5 43.5
-7 Video-ChatGPT 42.0
-8 Qwen-VL-Max 40.9
-9 Claude-3-Opus 40.2
-10 MA-LMM 38.9
-(e) Ego Reasoning
-Rank Model Score
-  Refer to caption GPT-4o 65.1
-Refer to caption GPT-4 Turbo 60.6
-Refer to caption InternVL-1.5 54.4
-4 Video-LLaVA 48.4
-5 Claude-3-Opus 47.8
-6 MiniGPT4-Video 44.5
-7 VideoChat2 43.8
-8 Qwen-VL-Max 43.3
-9 LLaVA 41.0
-10 MA-LMM 35.8
-(f) Plot QA
-Rank Model Score
-  Refer to caption GPT-4o 6.69
-Refer to caption GPT-4 Turbo 6.21
-Refer to caption Video-ChatGPT 5.48
-4 mPLUG-Owl-V 5.31
-5 Video-LLaVA 5.25
-6 Video-LLaMA-2 5.22
-7 LLaMA-VID 5.22
-8 Qwen-VL-Max 5.21
-9 VideoChat2 5.04
-10 VideoChat 5.01
-(g) Sub-Scene Caption.
-Rank Model Score
-  Refer to caption GPT-4o 56.7
-Refer to caption InternVL-1.5 32.8
-Refer to caption MovieChat 28.6
-4 GPT-4V 26.5
-5 LLaVA 25.5
-6 Video-ChatGPT 25.1
-7 MA-LMM 25.1
-8 Qwen-VL-Max 25.0
-9 TimeChat 24.7
-10 VideoChat 24.3
-(h) Action Order
-Rank Model Score
-  Refer to caption GPT-4o 46.3
-Refer to caption Video-LLaVA 35.9
-Refer to caption TimeChat 32.0
-4 Video-ChatGPT 31.1
-5 VideoChat2 29.6
-6 VideoChat 28.6
-7 LLaMA-VID 27.8
-8 Otter-V 26.7
-9 LLaVA 25.7
-10 Video-LLaMA-2 25.7
-(i) Action Count
-Rank Model Score
-  Refer to caption GPT-4o 64.6
-Refer to caption InternVL-1.5 50.4
-Refer to caption GPT-4 Turbo 49.2
-4 Video-LLaVA 47.3
-5 MiniGPT4-Video 44.5
-6 VideoChat2 44.5
-7 Qwen-VL-Max 42.2
-8 LLaVA-1.6 39.3
-9 Claude-3-Opus 36.5
-10 MA-LMM 36.4
-(j) M-Avg
-Rank Model Score
-  Refer to caption GPT-4o 5.80
-Refer to caption GPT-4 Turbo 5.35
-Refer to caption LLaMA-VID 4.22
-4 InternVL-1.5 4.02
-5 Qwen-VL-Max 3.96
-6 Video-ChatGPT 3.90
-7 mPLUG-Owl-V 3.84
-8 Video-LLaVA 3.84
-9 VideoChat2 3.81
-10 Video-LLaMA-2 3.78
-(k) G-Avg
-Table 6: Leaderboard of different tasks in MLVU.
-(until 2024/06/01)
-Appendix H More Experimental Results
-H.1 Leaderboard of each task in MLVU
-Table 6 shows the leaderboard of each task in our MLVU.
-H.2 Comprehensive evaluation results on MLVU Time-ladder
-Table 7, Table 8 and Table 9 show the experimental results on MLVU Time-ladder within 3 minutes, 6 minutes and 10 minutes.
-As anomaly recognition and topic reasoning are global tasks, thus, we can not annotate them in MLVU Time-ladder as the way provided in Appendix D.
-  Model Needle QA Ego Rea. Plot QA Action Or. Action Co. Sub-Scene Cap. Video Sum. M-Avg
-  Otter-I 25.0 25.0 25.4 25.0 25.0 4.04 2.59 25.1
-LLaVA-1.6 48.0 43.0 45.2 29.5 28.5 4.48 2.40 38.8
-InternVL-1.5 70.5 47.5 60.9 19.5 28.5 5.42 3.30 45.3
-Claude-3-Opus 34.5 42.5 38.1 15.5 22.0 4.12 3.35 30.5
-Qwen-VL-Max 49.5 42.2 50.3 27.7 16.3 5.58 2.95 37.2
-Otter-V 30.5 32.0 18.3 21.0 29.5 4.32 2.63 26.3
-mPLUG-Owl-V 31.0 29.5 22.3 29.0 24.0 5.35 2.68 27.2
-Video-LLaMA-2 55.0 35.0 36.5 22.5 22.0 5.76 2.47 34.3
-Video-ChatGPT 43.5 41.0 40.0 29.0 28.0 5.51 2.65 34.5
-VideoChat 32.5 31.5 26.4 24.5 26.0 5.39 2.61 28.2
-VideoChat2 54.0 53.5 50.3 30.0 31.0 5.11 3.25 43.8
-Video-LLaVA 58.0 47.5 55.8 29.5 34.5 5.38 2.89 45.0
-MiniGPT4-Video 57.0 44.5 47.2 27.5 26.5 4.38 2.98 40.5
-LLaMA-VID 38.5 36.5 36.0 22.5 25.5 5.31 3.55 31.8
-Movie-LLM 23.5 22.5 25.9 25.0 25.0 5.13 3.23 34.4
-MA-LMM 48.0 43.0 41.6 27.5 23.5 4.55 2.25 36.7
-MovieChat 32.5 29.5 27.4 21.0 21.0 3.60 2.60 26.3
-TimeChat 30.5 30.5 31.0 27.5 24.5 4.29 2.92 28.8
-Table 7: Quantitative results on MLVU Time-ladder (180s).“M-Avg” is calculated by the average of five multiple-choice tasks. Bold fonts mean the best results.
-  Model Needle QA Ego Rea. Plot QA Action Or. Action Co. Sub-Scene Cap. Video Sum. M-Avg
-  Otter-I 25.0 25.0 25.0 11.5 25.0 4.17 2.33 22.4
-LLaVA-1.6 41.0 38.5 39.6 22.0 26.0 4.14 2.18 33.4
-InternVL-1.5 66.5 48.5 58.4 34.0 24.0 5.23 3.12 46.2
-Claude-3-Opus 32.0 44.7 42.6 30.2 15.5 3.75 3.13 33.0
-Qwen-VL-Max 50.0 43.4 41.1 24.4 14.7 5.49 3.21 34.7
-Otter-V 27.0 30.5 23.9 13.0 26.5 4.34 2.51 24.2
-mPLUG-Owl-V 29.5 34.5 22.3 23.5 23.0 5.58 2.42 26.6
-Video-LLaMA-2 43.0 36.0 40.1 23.0 28.0 5.62 2.45 34.0
-Video-ChatGPT 42.5 35.0 27.4 23.5 32.0 5.22 2.56 32.1
-VideoChat 28.5 27.5 19.8 25.5 30.0 5.32 2.38 26.3
-VideoChat2 47.5 44.5 39.6 27.0 31.0 4.97 2.83 37.9
-Video-LLaVA 55.5 47.5 47.7 22.0 35.0 5.25 2.62 41.5
-MiniGPT4-Video 53.0 43.5 43.1 31.0 24.5 4.11 2.82 39.0
-LLaMA-VID 30.5 39.0 28.4 19.5 30.5 5.39 3.29 29.6
-Movie-LLM 26.0 22.5 32.0 22.5 26.0 5.16 3.14 25.8
-MA-LMM 41.5 37.0 41.6 27.0 27.5 4.52 2.19 34.9
-MovieChat 28.5 25.5 29.4 25.0 23.5 3.25 2.41 26.4
-TimeChat 36.5 30.0 31.0 26.5 26.0 4.37 2.66 30.0
-Table 8: Quantitative results on MLVU Time-ladder (360s).
-  Model Needle QA Ego Rea. Plot QA Action Or. Action Co. Sub-Scene Cap. M-Avg.
-  Otter-I 25.0 25.0 25.0 12.5 25.0 4.02 22.5
-LLaVA-1.6 37.5 38.5 43.9 25.0 28.5 4.45 34.6
-InternVL-1.5 56.0 46.0 55.6 33.0 25.0 5.17 43.1
-Claude-3-Opus 27.0 38.5 39.8 35.7 15.0 3.88 31.2
-Qwen-VL-Max 43.0 44.8 50.0 17.3 17.9 5.45 34.6
-Otter-V 27.5 28.5 22.4 14.0 26.0 4.09 23.7
-mPLUG-Owl-V 30.5 36.0 21.4 20.0 28.0 5.29 27.2
-Video-LLaMA-2 39.0 39.0 29.6 21.5 25.5 5.62 30.9
-Video-ChatGPT 41.0 40.5 28.6 24.5 32.5 5.37 33.4
-VideoChat 25.5 34.5 26.0 28.0 30.0 4.91 28.8
-VideoChat2 43.0 49.0 41.8 24.5 29.5 5.05 37.6
-Video-LLaVA 54.0 49.0 52.0 21.0 29.5 5.21 41.1
-MiniGPT4-Video 54.0 39.0 41.8 25.0 28.0 4.13 37.6
-LLaMA-VID 32.0 36.5 34.2 27.0 27.0 5.39 31.3
-Movie-LLM 26.5 32.5 27.0 21.5 24.5 5.25 26.4
-MA-LMM 44.5 41.5 35.7 24.0 28.0 4.72 34.7
-MovieChat 27.5 32.5 22.4 24.0 27.0 3.48 26.7
-TimeChat 28.0 32.0 23.5 26.0 24.5 4.19 26.8
-Table 9: Quantitative results on MLVU Time-ladder (600s).
 Appendix I More Visualized Examples of MLVU.
+We present additional visualizations of our MLVU annotation examples in Figures 10, 11, and 12.
 Figure 10: More Examples of Topic Reasoning and Video Summarization Tasks.
@@
 Figure 12: More Examples of Sub-Scene Captioning.
-Appendix J Licensing, Hosting and Maintenance Plan
-Author Statement.
-We bear all responsibilities for the licensing, distribution, and maintenance of our dataset.
-License.
-MLVU is under CC-BY-NC-SA 4.0 license.
-Hosting.
-MLVU can be viewed and downloaded on GitHub at https://github.com/JUNJIE99/MLVU or on Huggingface at https://huggingface.co/datasets/MLVU/MVLU.
-We assure its long-term preservation for future reference and use.
-The annotations for questions and answers are provided in the JSON file format, while the raw videos are available in the MP4 format.
-For the video files, we do not hold any copyright, and all video rights belong to the video authors.
-To facilitate user usage, under the premise of a user agreement that the data will be used solely for research purposes and not for commercial use, we provide the download methods for these videos.
-We have taken measures such as reducing resolution, modifying aspect ratio, and editing to minimize the impact on the original work rights.
-However, there is still a risk that copyright holders may request the removal of some data.
-If this happens, we will follow the practice of Movienet [13] to sparsely collect these video frames.
-This will not significantly impact our data usage, as all our annotations are focused on the visual information in the videos, unrelated to the audio.
-Moreover, most of the models we evaluate are essentially still frame-based video processing and do not involve audio processing.
-If retaining video frames is also not allowed, we will still preserve the annotation data and provide metadata for the corresponding videos.
-Metadata.
-Appendix K Datasheet
-K.1 Motivation
-For what purpose was the dataset created?
-Answer: The creation of MLVU is to facilitate the evaluation and development of MLLM’s long video understanding capabilities.
-Compared to previous video understanding benchmarks, MLVU possesses a substantial extension of video length, diversified video categories, and diversified evaluation tasks.
-MLVU is the first comprehensive benchmark for long video understanding.
-Who created the dataset (e.g., which team, research group) and on behalf of which entity (e.g., company, institution, organization)?
-Answer: The MLVU is created by Junjie Zhou (BUPT, BAAI), Yan Shu (BAAI), Bo Zhao (BAAI), Boya Wu (BAAI), Shitao Xiao (BAAI), Xi Yang (BAAI), Yongping Xiong (BUPT), Bo Zhang (ZJU), Tiejun Huang (PKU, BAAI), and Zheng Liu (BAAI)
-Who funded the creation of the dataset?
-Answer: Beijing Academy of Artificial Intelligence, Beijing University of Posts and Telecommunications, Peking University, and Zhejiang University.
-K.2 Composition
-What do the instances that comprise the dataset represent? (e.g., documents, photos, people, countries)
-Answer: Each instance in our dataset represents a long video ranging from 3 minutes to 2 hours in duration, a question, a standard answer, and for the multiple-choice task, there are also 4 options.
-Videos are stored in MP4 file format, while the questions, standard answers, and options are all stored in JSON format files.
-How many instances are there in total (of each type, if appropriate)?
-Answer: In total, we collect 2593 instances, including 2175 multiple-choice questions and 418 free-form generation questions.
-The number of distinct videos is 1334, as one video may correspond to multiple questions.
-The specific distribution of question types and video duration can be found in Figure 1 in the main paper.
-Does the dataset contain all possible instances or is it a sample (not necessarily random) of instances from a larger set?
-Answer: Only a part of the questions was sampled from existing datasets (the ego reasoning task and the anomaly recognition task).
-The rest of the questions were newly annotated by us (for example, plot question answering, sub-scene captioning, video summarization, topic reasoning), or they were substantially modified based on existing annotations (needle question answering, action count, action order).
-During the data collection and annotation process, we ensured that the instances within each task were diverse in terms of videos and questions.
-Is there a label or target associated with each instance?
-Answer: Yes, for the multiple-choice question type, each instance provides the correct option. For the free-form generation task, each instance comes with a standard caption or summary answer.
-Is any information missing from individual instances?
-Answer: All instances are complete.
-Are relationships between individual instances made explicit (e.g., users’ movie ratings, social network links)?
-Answer: Some instances may have the same video but different questions and answers. Each instance clearly indicates the corresponding video, and each video has a unique identifier.
-Are there recommended data splits (e.g., training, development/validation, testing)?
-Answer: No. Our MLVU is specifically designed for evaluation, with its core objective being to assess the capability of MLLMs to understand long-term video.
-Are there any errors, sources of noise, or redundancies in the dataset?
-Answer: No.
-Is the dataset self-contained, or does it link to or otherwise rely on external resources (e.g., websites, tweets, other datasets)?
-Answer: All data will be publicly accessible in the dataset repository.
-Our annotations will be stored in JSON format.
-As our collected videos include parts of movies, TV series, documentaries, and cartoons, we have taken measures such as reducing resolution, modifying aspect ratio, and editing to minimize the impact on the original work rights.
-However, there is still a risk that copyright holders may request removal of some data.
-If this happens, we will follow the practice of Movienet [13] to sparsely collect these video frames.
-This will not have much impact on our data usage, as all our annotations are focused on the visual information in the videos, unrelated to the audio.
-If retaining video frames is also not allowed, we will still keep the annotation data and provide metadata for the corresponding videos.
-Does the dataset contain data that might be considered confidential?
-Answer: No.
-Does the dataset contain data that, if viewed directly, might be offensive, insulting, threatening, or might otherwise cause anxiety?
-Answer: No.
-K.3 Collection Process
-The data collection process is described in Section 3.2 of the main paper and Appendix B, D, and E.
-K.4 Uses
-Has the dataset been used for any tasks already?
-Answer: Yes, MLVU has been used to evaluate the long-video understanding capabilities of as many as 20 different MLLMs. For specific details, please refer to Section 4.
-What (other) tasks could the dataset be used for?
-Answer: MLVU is primarily used to comprehensively evaluate the long-video understanding capabilities of MLLMs.
-Since MLVU includes multiple different tasks, it can also be used to individually assess the ability of MLLMs or other video-specific models on a particular task, such as video summarization.
-Is there a repository that links to any or all papers or systems that use the dataset?
-Answer: No.
-Is there anything about the composition of the dataset or the way it was collected and preprocessed/cleaned/labeled that might impact future uses?
-Answer: As our collected videos include parts of movies, TV series, documentaries, and cartoons, we have taken measures such as reducing resolution, modifying aspect ratio, and editing to minimize the impact on the original work rights.
-However, there is still a risk that copyright holders may request removal of some data.
-If this happens, we will follow the practice of Movienet [13] to sparsely collect these video frames.
-This will not have much impact on our data usage, as all our annotations are focused on the visual information in the videos, unrelated to the audio.
-If retaining video frames is also not allowed, we will still keep the annotation data and provide metadata for the corresponding videos.
-Are there tasks for which the dataset should not be used?
-Answer: The MLVU cannot be used to evaluate the image or video generation capabilities of the MLLM.
-K.5 Distribution
-Will the dataset be distributed to third parties outside of the entity (e.g., company, institution, organization) on behalf of which the dataset was created?
-Answer: Yes. The benchmark is publicly available on the Internet.
-How will the dataset will be distributed (e.g., tarball on website, API, GitHub)?
-Answer: The benchmark is available on GitHub at https://github.com/JUNJIE99/MLVU or on Huggingface at https://huggingface.co/datasets/MLVU/MVLU.
-Will the dataset be distributed under a copyright or other intellectual property (IP) license, and/or under applicable terms of use (ToU)?
-Answer: CC-BY-NC-SA 4.0.
-Have any third parties imposed IP-based or other restrictions on the data associated with the instances?
-Answer: No.
-Do any export controls or other regulatory restrictions apply to the dataset or to individual instances?
-Answer: No.
-K.6 Maintenance
-Who will be supporting/hosting/maintaining the dataset?
-Answer: The authors will be supporting, hosting, and maintaining the dataset.
-How can the owner/curator/manager of the dataset be contacted (e.g., email address)?
-Answer: Please contact the official email of our project (mlvubenchmark@gmail.com), or contact the one of the authors (zhoujunjie@bupt.edu.cn; shuyan9812@gmail.com)
-Is there an erratum?
-Answer: No. We will make announcements if there are any.
-Will the dataset be updated (e.g., to correct labeling errors, add new instances, delete instances)?
-Answer: Yes. We will post new update in https://github.com/JUNJIE99/MLVU and https://huggingface.co/datasets/MLVU/MVLU if there is any.
-If the dataset relates to people, are there applicable limits on the retention of the data associated with the instances (e.g., were individuals in question told that their data would be retained for a fixed period of time and then deleted)?
-Answer: People may appear in the reference videos. People may contact us to exclude specific data instances if they appear in the reference videos.
-Will older versions of the dataset continue to be supported/hosted/maintained?
-Answer: Yes. Old versions will also be hosted in https://github.com/JUNJIE99/MLVU and https://huggingface.co/datasets/MLVU/MVLU.
-If others want to extend/augment/build on/contribute to the dataset, is there a mechanism for them to do so?
-Answer: If others wish to add data, they can apply to do so provided the data is compliant and reasonable. However, making other modifications based on our dataset is currently not allowed.